### v1

# Bulk-Boundary Correspondence and Exceptional Points for a Dimerized Hatano-Nelson Model with Staggered Potentials

Abstract
It is well-known that the standard bulk-boundary correspondence does not hold for non-Hermitian systems in which also new phenomena such as exceptional points do occur. Here we study by analytical and numerical means a paradigmatic one-dimensional non-Hermitian model with dimerization, asymmetric hopping, and imaginary staggered potentials. We present analytical solutions for the eigenspectrum of this model with both open and closed boundary conditions as well as for the singular-value spectrum. We explicitly demonstrate the proper bulk-boundary correspondence between topological winding numbers in the periodic case and singular values in the open case. We also show that a non-trivial topology leads to protected eigenvalues in the entanglement spectrum. In the $\mathcal{PT}$-symmetric case, we find that the model has a phase where exceptional points become dense in the thermodynamic limit.
I Introduction
It is well known that fundamental principles in quantum mechanics are formulated using Hermitian operators, which assure for example unitary time evolution of closed systems. However, practically, no quantum system is perfectly isolated. The way a system interacts with its environment is often represented through a Master equation like the Lindblad equation [1]. If quantum jumps can be neglected, then the time evolution described by a Lindblad equation can be replaced by an effective non-Hermitian (NH) Hamiltonian [2]. The validity of this effective approach has been verified in different experimental arrangements [3, 4, 5]. Strictly speaking, it requires a post-selection of those time evolutions which stay in the considered manifold [6].
Non-Hermitian systems exhibit distinct properties associated with their complex spectrum and the difference between left and right eigenvectors [7]. In this context, as with Hermitian systems, Gaussian non-Hermitian models can offer valuable insights into possible phases and phenomena. In particular, the topological properties of Gaussian non-Hermitian models have been studied and new phenomena have been revealed [8, 9, 10, 11, 12, 13, 11, 14, 15, 16]. For example, non-hermiticity leads to a proliferation of possible symmetry classes: the famous Altland-Zirnbauer ten-fold classification [17] for Hermitian Gaussian models is replaced by a 38-fold classification [8, 14] in the non-Hermitian domain. The richer classification scheme is a direct consequence of the distinction between transposition and conjugate transposition. Other distinctive properties include the extreme sensitivity of the eigenspectrum to the boundary conditions [15] leading to the skin effect and the breakdown of the conventional bulk-boundary correspondence [10, 12, 18, 19]. Another remarkable aspect of non-Hermitian Hamiltonians is the existence of so-called exceptional points [20] where both eigenvalues and eigenvectors coalesce. Such exceptional points can be detected in experiments [6].
In order to highlight the different phenomena possible in Gaussian non-Hermitian models, we will study a minimal model which allows to turn various symmetries on and off. In the Hermitian case, one of the best-known one-dimensional systems with symmetry-protected topological order is the Su-Schrieffer-Heeger (SSH) model [21] which has chiral symmetry and belongs to the BDI class. This model has alternating strong and weak hoppings and topological edge modes are present if the two sites at the end of an open chain are weakly coupled to their respective neighbors. Here we make the model non-Hermitian by allowing for parity breaking hoppings, i.e., different hopping amplitudes to the left and to the right. In addition, we also introduce staggered imaginary onsite potentials to allow for a non-Hermitian, $\mathcal{PT}$-symmetric limit. Physically, the parity breaking can arise in an effective description of an SSH model with gain and loss. This phenomenon has been investigated in plasmonic chains [22] and photonic lattices [23]. For equal hoppings to the left and to the right, the model is $\mathcal{PT}$ symmetric. The model is depicted in Fig. 1
[…]
Figure 1: The considered minimal model has a two-site unit cell, asymmetric intra- ($V_{L},V_{R}$) and inter-cell ($W_{L},W_{R}$) hopping as well as complex staggered potentials ±$\pm iu$.
and can be viewed either as a dimerized Hatano-Nelson model [24] or an SSH chain with asymmetric hopping.
Our paper is organized as follows: In Sec. II, we introduce the model and the difference equations the right and left eigenvectors have to fulfill. In Sec. III, we diagonalize the model for periodic and aperiodic boundary conditions. Furthermore, we determine the parameters for which the excitation gap closes and for which exceptional points do occur. We also introduce and determine the two winding numbers for this system. In Sec. IV, we then analytically study the open boundary case. As expected, the spectrum and the eigenvectors change completely as compared to the closed boundary case. This is also true for the conditions for the closing of the excitation gap and for the location of exceptional points. In Sec. V, we present some of the central results of our paper. We analytically calculate the singular value spectrum of a finite open chain as well as the zero-energy eigenstates of a semi-infinite chain. Based on these results we explicitly establish a bulk-boundary correspondence for this specific model following results recently proven in general in Ref. [25]. In Sec. VI, we then discuss the entanglement entropy when cutting the system in half. The studied model is Gaussian and therefore its correlation matrix can be obtained [26, 27, 28, 29, 30]. We find that in topological phases there are topologically protected eigenvalues in the entanglement spectrum. The last section is devoted to a short summary and our conclusions.
II The dimerized Hatano-Nelson model
We consider a model with a two-site unit cell with asymmetric hopping in both intra ($V_{L},V_{R}$) and inter ($W_{L},W_{R}$) cells, see Fig. 1. Complex staggered fields ±$\pm\text{i}u$ act on the sites of the chain. This model can be viewed either as a dimerized Hatano-Nelson model or as a Su-Schrieffer-Heeger (SSH) model with asymmetric hopping. It is described by the Hamiltonian
$\mathcal{H}$ $=$ $\sum_{j=1}^{\lfloor\frac{N-1}{2}\rfloor}W_{L}c_{2j}^{\dagger}c_{2j+1}+W_{R}c_{2j+1}^{\dagger}c_{2j}$ (II.1)
$+$ $\gamma(W_{R}c_{1}^{\dagger}c_{L}+W_{L}c_{L}^{\dagger}c_{1})$
$+$ $\sum_{j=1}^{\lfloor\frac{N}{2}\rfloor}V_{L}c_{2j-1}^{\dagger}c_{2j}+V_{R}c_{2j}^{\dagger}c_{2j-1}$
-- $\text{i}u\sum_{j=1}^{\lfloor\frac{N}{2}\rfloor}c_{2j}^{\dagger}c_{2j}+\text{i}u\sum_{j=1}^{\lfloor\frac{N+1}{2}\rfloor}c_{2j-1}^{\dagger}c_{2j-1},$
where $c_{j}$ and $c_{j}^{\dagger}$ are fermionic operators satisfying the algebra
$\{c_{j},c_{k}\}=\{c_{j}^{\dagger},c_{k}^{\dagger}\}=0,\quad\{c_{j},c_{k}^{\dagger}\}=\delta_{j,k}$ (II.2)
and $\lfloor x\rfloor$ denotes the floor function. By writing the Hamiltonian in the form (II.1), we can consider different boundary conditions (controlled by the parameter $\gamma$) in a uniform fashion. In most parts of the paper, we assume that the parameters $W_{L,R}$, $V_{L,R}$ and $u$ are real. However, when discussing the exceptional points of this model it makes sense to consider complex parameters as well.
The Hamiltonian (II.1) can be written in the compact form
$\mathcal{H}=\bm{c}^{\dagger}\mathcal{T}_{N}\bm{c}$ (II.3)
where $\mathcal{T}_{N}$ is an asymmetric (almost) tridiagonal matrix with dimension $N\times N$ with matrix elements
$t_{2j-1,2j}=V_{L},\quad t_{2j,2j-1}=V_{R},\quad t_{2j,2j}=-\text{i}u,$
$t_{2j,2j+1}=W_{L},\quad t_{2j+1,2j}=W_{R},\quad t_{2j-1,2j-1}=\text{i}u,$
$t_{1,N}=\gamma W_{R},\quad t_{N,1}=\gamma W_{L},$ (II.4)
and $\bm{c}^{\dagger}=(c_{1}^{\dagger},\dots,c_{N}^{\dagger})$, $\bm{c}=(c_{1},\dots,c_{N})^{T}$ with $T$ denoting transposition.
To make this paper self-contained, we will recall the diagonalization of the asymmetric SSH chain for both closed ($\gamma=\pm 1)$ and open ($\gamma=0$) boundary conditions, referred to respectively as CBC or OBC. In both cases, the diagonalization of $\mathcal{T}_{N}$ with entries (II) follows from difference equations of the spectral problem. Here the lack of hermiticity of the single body Hamiltonian implies that left and right eigenvectors have to be considered. The spectral problems read
$\mathcal{T}_{N}\vec{r}=\epsilon~{}\vec{r},\quad(\mathcal{T}_{N})^{T}\vec{\ell}=\epsilon~{}\vec{\ell},$ (II.5)
where $\vec{x}=(x_{1},\dots,x_{N})^{T}$ are right ($x=r$) and left ($x=\ell$) eigenvectors of $\mathcal{T}_{N}$, while $\epsilon$ is an eigenvalue. The components of the right eigenvector satisfy the coupled difference equations
$V_{L}r_{2j}+W_{R}r_{2j-2}=(\epsilon-\text{i}u)r_{2j-1},$ (II.6)
$W_{L}r_{2j+1}+V_{R}r_{2j-1}=(\epsilon+\text{i}u)r_{2j},$ (II.7)
with the boundary conditions
$r_{0}=\gamma r_{N},\quad r_{N+1}=\gamma r_{1}.$ (II.8)
Exchanging $R\leftrightarrow L$ one obtains the analogous equations for $\ell_{j}$, namely
$V_{R}\ell_{2j}+W_{L}\ell_{2j-2}=(\epsilon-\text{i}u)\ell_{2j-1},$ (II.9)
$W_{R}\ell_{2j+1}+V_{L}\ell_{2j-1}=(\epsilon+\text{i}u)\ell_{2j},$ (II.10)
with the boundary conditions,
$\ell_{0}=\gamma\ell_{N},\quad\ell_{N+1}=\gamma\ell_{1}.$ (II.11)
For $\gamma=0$, we consider (II.8, II.11) for both parities of $N$, while for $\gamma\neq 0$ we only consider the case where $N$ is even.
In the following sections, each boundary condition is considered separately.
III Closed boundary conditions
III.1 Diagonalization
Let us briefly recapitulate the diagonalization of $\mathcal{T}_{N}$ for closed boundary conditions (CBC), that is, $\gamma=\pm 1$. In this case, $\mathcal{T}_{N}$ is a block circulant matrix and can be easily diagonalized. We only consider $N$ even for CBC. The difference equations (II.6,II.7,II.8) and (II.9,II.10,II.11) are easily solved by the ansatz
$r_{2j-1}=a_{R}(k)e^{-\text{i}kj},\quad r_{2j}=b_{R}(k)e^{-\text{i}kj},$ (III.1)
$\ell_{2j}=a_{L}(k)e^{\text{i}kj},\quad\ell_{2j-1}=b_{L}(k)e^{\text{i}kj}$ (III.2)
where $k$ is a free parameter. The parameters $a_{R,L}(k)$ and $b_{R,L}(k)$ are fixed by the spectral problem of the two band Bloch Hamiltonian
$H(k)=\left(\begin{array}[]{cc}\text{i}u&H_{1}(k)\\ H_{2}(k)&-\text{i}u\\ \end{array}\right),$ (III.3)
where
$H_{1}(k)$ $=$ $V_{L}+W_{R}e^{\text{i}k},$ (III.4)
$H_{2}(k)$ $=$ $V_{R}+W_{L}e^{-\text{i}k}.$ (III.5)
That is
$H(k)|r^{\pm}(k)\rangle=\epsilon^{\pm}(k)|r^{\pm}(k)\rangle,$
±$\langle\ell^{\pm}(k)|H(k)=\epsilon^{\pm}(k)\langle\ell^{\pm}(k)|,$ (III.6)
where the quasienergies are given by
$\epsilon^{\pm}(k)=\pm\sqrt{H_{1}(k)H_{2}(k)-u^{2}},$ (III.7)
and the left and right eigenvectors by
$|r^{\pm}(k)\rangle$ $=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{c}a_{R}^{\pm}(k)\\ b_{R}^{\pm}(k)\\ \end{array}\right)$ (III.10)
$=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{c}\text{i}u\pm\sqrt{H_{1}(k)H_{2}(k)-u^{2}}\\ H_{2}(k)\\ \end{array}\right),$ (III.13)
±$\langle\ell^{\pm}(k)|$ $=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{cc}a_{L}^{\pm}(k),&b_{L}^{\pm}(k)\end{array}\right),$ (III.15)
$=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{cc}\text{i}u\pm\sqrt{H_{1}(k)H_{2}(k)-u^{2}},&H_{1}(k)\,.\end{array}\right)$ (III.17)
Normalizing left and right eigenvectors according to the biorthogonality condition
$\langle\ell^{\epsilon}(k)|r^{\epsilon^{\prime}}(k)\rangle=\delta_{\epsilon,\epsilon^{\prime}},$ (III.19)
fixes the normalization constant
$N^{\pm}=2\left(H_{1}(k)H_{2}(k)-u^{2}\pm\text{i}u\sqrt{H_{1}(k)H_{2}(k)-u^{2}}\right).$ (III.20)
To conclude the solution, the parameter $k$ is quantized by (II.8) and given by
$k_{m}^{\text{(PBC)}}=\frac{2\pi(2m)}{N},\quad m=1,\dots,\frac{N}{2},$ (III.21)
for periodic boundaries and
$k_{m}^{\text{(APBC)}}=\frac{2\pi(2m+1)}{N},\quad m=1,\dots,\frac{N}{2},$ (III.22)
for antiperiodic boundaries.
III.2 Gap closing and exceptional points
The bands (III.7) for the asymmetric SSH chain are in general complex, and a band touching requires the vanishing of both the real and the imaginary parts of the complex gap
$\Delta_{c}=(\epsilon^{+}(k)-\epsilon^{-}(k))/2=\sqrt{H_{1}(k)H_{2}(k)-u^{2}}.$ (III.23)
The closing of both the real and the imaginary parts is well captured by the non-Hermitian gap [31]
$\Delta=\text{min}_{k}|\Delta_{c}(k)|.$ (III.24)
The complex gap (III.23) closes when,
$H_{1}(k)H_{2}(k)-u^{2}=0.$ (III.25)
For each value of $k$ quantized by (III.21,III.22), the condition (III.25) leads to constraints in the model parameters and can determine a rich variety of phase boundaries. Furthermore, it follows immediately from Eqs. (III.13,III.17) that at points where the complex gap (III.25) vanishes the eigenvectors $\{|r^{+}(k)\rangle,|r^{-}(k)\rangle\}$ and also $\{\langle\ell^{+}(k)|,\langle\ell^{-}(k)|\}$ coalesce. Gap closing points are therefore also always exceptional points. Exceptional points and their relation with the discriminant of the characteristic polynomial of the hopping matrix $\mathcal{T}_{N}$ are discussed further in App. B.
To analyze the manifold of exceptional points given by Eq. (III.25) it is convenient to first set $u=0$. In this case, the system has sublattice symmetry $\sigma^{z}H(k)\sigma^{z}=-H(k)$. [^1] In non-Hermitian systems, chiral and sublattice symmetries are not equivalent [14]. Indeed for $u=0$ the chiral symmetry $\sigma^{z}H(k)^{\dagger}\sigma^{z}=-H(k)$ is not satisfied since $\sigma^{z}H(k)^{\dagger}\sigma^{z}=-H(k)^{\dagger}$. Then, for every allowed quantized $k_{m}$ value and relaxing the requirement of real couplings, the solutions of $H_{1}=0$ or $H_{2}=0$ are
$\left(\frac{V_{L}}{W_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}\quad\text{or}\quad\left(\frac{W_{L}}{V_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}.$ (III.26)
where $k_{m}$ is given by (III.21,III.22). It follows that for most values of $k_{m}$ the coupling parameters which lead to a gap closing/exceptional point are complex. For finite $N$ and periodic boundaries, one can see that $m=N/2$ ($k=2\pi$) and $m=N/4$ ($k=\pi$, if $N/2$ is even) are the only possible values of $k$ that leads to real ratios,
$\left(\frac{V_{L}}{W_{R}}\right)_{\text{EP}}=\pm 1\quad\text{or}\quad\left(\frac{W_{L}}{V_{R}}\right)_{\text{EP}}=\pm 1.$ (III.27)
Fixing $V_{R}=W_{R}=1$ (no loss in generality if $u=0$), these exceptional lines can be visualized as a function of $V_{L}$ and $W_{L}$, see the left panel in Fig. 2. For finite $N$ and antiperiodic boundaries, the ratios (III.26) are real for $m=1/2(N/2-1)$ ($k=\pi$, if $N/2-1$ is even). In the thermodynamic limit, the unit circles are filled and real ratios occur at $k=\pi$ or $k=2\pi$ for both periodic and antiperiodic boundary conditions. We remark that the sublattice-symmetric model has been widely considered in the literature, see e.g. Refs. [11, 13, 32, 33, 34, 35]. Here we emphasize that the gap closing points and exceptional points coincide and that they can also occur in the complex parameter plane for every allowed, quantized value $k_{m}$.
Figure 2: The non-Hermitian gap $\Delta$ as a function of the model parameters for $N=100$ and periodic boundary conditions. In panels (a) and (b) we set $V_{R}=W_{R}=1$ and $u=0$ and $u=0.5$ respectively. In panel (c) we consider $V_{L}=V_{R}=V$, $W_{L}=W_{R}=W$ and the field $u=1$. The numbers in panel (a) indicate the two winding numbers $(\nu_{1},\nu_{2})$, the numbers in panel (b) the winding number $I_{1}$.
Now we consider a nonzero real field $u$. The field couples $H_{1}(k)$ and $H_{2}(k)$ and Eq. (III.25) no longer factorizes, giving a constraint for all model parameters for every possible value of $k$. Eq. (III.25) can be solved for any of the model parameters, leading in general to a complex number. As an example, in the central panel of Fig. 2, the non-Hermitian gap (III.24) is shown for fixed parameters $V_{R}=W_{R}=1$ and $u=0.5$ as a function of $V_{L}$ and $W_{L}$. In this case, we can observe the merging of the zero field phases $(\nu_{1},\nu_{2})=(0,0)$ and $(1,-1)$ (we discuss the meaning of these numbers below) into a single phase as well as the bending of the phases boundaries.
Another interesting case to consider is the $\mathcal{PT}$ symmetric case with $V_{L}=V_{R}=V$ and $W_{L}=W_{R}=W$ but the real field $u$ is arbitrary, such that $\sigma^{x}H(k)\sigma^{x}=H(k)^{*}$. In this case, the constraint (III.25) reduces to,
$V^{2}+W^{2}+2VW\cos(k_{m})-u^{2}=0\,.$ (III.28)
To visualize the gap closing, we can set $u=1$ (no loss of generality) and plot the gap (III.24) as a function of $V$ and $W$, see the right panel of Fig. 2. The dark blue lines mark the the gap closing/ exceptional points, which only occur within the regions $|V-W|<1<|V+W|$ or $|V+W|<1<|V-W|$. In the thermodynamic limit, this region is gapless and filled with exceptional points. It thus constitutes an entire exceptional phase.
We end this subsection remarking that in addition to the collision of quasienergies $\Delta_{c}=0$, another type of degeneracy in the spectrum can also be found. This degeneracy occurs within each band ϵ±superscriptitalic-$\epsilon^{\pm}$ for different quasimomentum $k$. Namely, one can impose
$H_{1}(k)H_{2}(k)=H_{1}(k^{\prime})H_{2}(k^{\prime})$ (III.29)
which gives another constraint on the model parameters. However, these degeneracies do not lead to exceptional points, as the eigenvectors do not coalesce, recall (III.13,III.17). This type of degeneracy, however, does appear as a factor in the discriminant of the characteristic polynomial of $\mathcal{T}_{N}$, see further remarks on this point in App. B, in particular Eq. (B.3).
III.3 Winding numbers
The eigenvalues of a non-Hermitian Hamiltonian are in general complex, and can trace a closed path around an arbitrary reference complex energy $E_{B}$, in which case a winding number can be assigned [36, 37, 18, 38, 32],
[…]
provided that $H(k)$ has a point gap, that is, $\det(H(k)-E_{B})\neq 0$. For concreteness we will chose $E_{B}=0$. Observe that the definition (III.30) is universal since no symmetry requirements are made. Therefore, one dimensional non-Hermitian systems—in contrast to Hermitian ones—can have non-trivial topology in the absence of any spatial or non-spatial symmetries.
If symmetries are present, then they lead to additional topological invariants [14]. So far, mostly the case of a two-band model with sub-lattice symmetry has been considered [39, 40, 41, 42, 11, 38]. The Hamiltonian of such a model takes the form
$H(k)=\left(\begin{array}[]{cc}0&\hat{H}_{1}(k)\\ \hat{H}_{2}(k)&0\\ \end{array}\right)\,.$ (III.31)
Using the polar decomposition $\hat{H}_{j}(k)=d_{j}(k)q_{j}(k)$ with $q_{j}(k)$ unitary and $d_{j}(k)$ positive semi-definite, one can define two winding numbers
[…]
Instead of the winding numbers $\nu_{1,2}$, one can equivalently consider [13, 43]
$I_{1}=\nu_{1}+\nu_{2},$ (III.33)
which is consistent with the definition of $I_{1}$ in Eq. (III.30), and
$I_{2}=\nu_{2}-\nu_{1}$ (III.34)
as independent winding numbers. For our model with $u=0$, we have $H_{j}(k)=|H_{j}(k)|e^{\text{i}\phi_{j}}$ and evaluating the integrals (III.33,III.34) leads to
$I_{1}(u=0)=\begin{cases}-1,&\text{if}\quad~{}\left|\frac{W_{L}}{V_{R}}\right|>1\land\left|\frac{V_{L}}{W_{R}}\right|>1\\ 1,&\text{if}\quad\left|\frac{W_{L}}{V_{R}}\right|<1\land\left|\frac{V_{L}}{W_{R}}\right|<1\\ 0,&\text{otherwise}\end{cases}$ (III.35)
and
$I_{2}=\begin{cases}-2&\text{if}\quad~{}\left|\frac{V_{L}}{W_{R}}\right|<1\land\left|\frac{W_{L}}{V_{R}}\right|>1\\ 0&\text{if}\quad\left|\frac{V_{L}}{W_{R}}\right|>1\land\left|\frac{W_{L}}{V_{R}}\right|<1\\ -1&\text{otherwise}.\end{cases}$ (III.36)
Clearly, the winding numbers are not defined when $\left|\frac{V_{L}}{W_{R}}\right|=\left|\frac{W_{L}}{V_{R}}\right|=1$ which are precisely the gapless/exceptional lines discussed in the previous subsection. The two winding numbers $\nu_{1,2}$ are shown in the left panel of Fig. 2 for parameters $V_{R}=W_{R}=1$ (no loss of generality) in each phase as a function of $V_{L}$ and $W_{L}$.
Next, we briefly discuss the $\mathcal{PT}$-symmetric case. The Hamiltonian (III.3) for $V_{L}=V_{R}=V$ and $W_{L}=W_{R}=W$ satisfies the $\mathcal{PT}$ symmetry $\sigma^{x}H(k)\sigma^{x}=H(k)^{*}$ as well as chiral symmetry $\sigma^{z}H(k)^{\dagger}\sigma^{z}=-H(k)$, and this case is part of the AZ symmetry class AIII [14]. The phase diagram for $u=1$ is shown in Fig. 2(c) and consists of gapped phases and phases where exceptional points become dense in the thermodynamic limit. These phases will be discussed further in Sec. VI when we analyze the entanglement spectrum.
IV Open boundary conditions
IV.1 Diagonalization
Let us also recall the diagonalization of $\mathcal{T}_{N}$ (II) for open boundary conditions (OBC), that is, $\gamma=0$. In this case, $\mathcal{T}_{N}$ is an asymmetric tridiagonal 2-Toeplitz matrix whose eigenvalue problem was solved from different perspectives in Refs. [44, 45]. In App. A, we provide some details on both approaches. The connection between these two approaches was pointed out in Ref. [46] and the method [45] was previously applied to the Hermitian SSH chain in Refs. [47, 48].
The solution of the open boundary case can also be obtained by a simple transformation to the standard (symmetric) SSH chain [12], namely,
$S\mathcal{T}_{N}S^{-1}=\mathcal{T}_{N}^{SSH}$ (IV.1)
where $S$ is a diagonal matrix with entries
$s_{2j-1}=\frac{V_{R}^{\frac{j-1}{2}}W_{R}^{\frac{j-1}{2}}}{V_{L}^{\frac{j-1}{2}}W_{L}^{\frac{j-1}{2}}},\quad s_{2j}=\frac{V_{R}^{\frac{j}{2}}W_{R}^{\frac{j}{2}}}{V_{L}^{\frac{j}{2}}W_{L}^{\frac{j}{2}}}\frac{W_{L}}{W_{R}},$ (IV.2)
and $\mathcal{T}_{N}^{SSH}$ is a symmetric hopping matrix of the standard SSH chain with effective couplings,
$V_{\text{eff}}=\sqrt{V_{L}V_{R}},\quad W_{\text{eff}}=\sqrt{W_{L}W_{R}}.$ (IV.3)
The quasienergies are then given by
$\epsilon^{\pm}=\pm\sqrt{\tilde{H}_{1}(\theta)\tilde{H}_{2}(\theta)-u^{2}},$ (IV.4)
where
$\tilde{H}_{1}(\theta)=V_{L}(e^{i\theta/2}+e^{-i\theta/2}\delta^{-1/2}),$ (IV.5)
$\tilde{H}_{2}(\theta)=V_{R}(e^{-i\theta/2}+e^{i\theta/2}\delta^{-1/2}),$ (IV.6)
with
$\sqrt{\delta}=\frac{\sqrt{V_{L}}\sqrt{V_{R}}}{\sqrt{W_{L}}\sqrt{W_{R}}},$ (IV.7)
and $\theta$ is a parameter that is given by different expressions depending on the parity of $N$. We remark that, for odd $N$, there is an isolated root $\epsilon=\text{i}u$ which is not contained in (IV.4).
For odd $N$, the parameter $\theta$ is simply
$\theta_{m}^{(OBC)}=\frac{2\pi m}{N+1},\quad m=1,\dots,\frac{N-1}{2},$ (IV.8)
which gives $N-1$ eigenvalues from (IV.4) and (IV.8). Together with the isolated root $\epsilon=\text{i}u$, one has all the $N$ eigenvalues of the tridiagonal matrix when $\gamma=0$.
For even $N$, the parameter $\theta$ is determined via the transcendental equation
$\sqrt{\delta}\sin\left(\left(\frac{N}{2}+1\right)\theta\right)+\sin\left(\frac{N}{2}\theta\right)=0\,.$ (IV.9)
We remark that the transcendental equation (IV.9) has the same form as the transcendental equation of the quantum Ising chain with transverse field $\lambda=\sqrt{\delta}$ and $L/2$ sites [49]. Meanwhile, the eigenvector components can be written as
$r_{2n}$ $=$ $\frac{V_{R}^{\frac{n-1}{2}}W_{R}^{\frac{n-1}{2}}}{V_{L}^{\frac{n-1}{2}}W_{L}^{\frac{n-1}{2}}}(\epsilon-\text{i}u)\frac{\sin(n\theta)}{\sin(\theta)},$ (IV.10)
$r_{2n-1}$ $=$ $\frac{V_{R}^{\frac{n-1}{2}}W_{R}^{\frac{n-1}{2}}}{V_{L}^{\frac{n-1}{2}}W_{L}^{\frac{n-1}{2}}}\frac{\left(\sin(n\theta)+\delta^{-\frac{1}{2}}\sin((n-1)\theta)\right)}{V_{L}^{-1}\sin(\theta)},$
and
$\ell_{2n}$ $=$ $\frac{V_{L}^{\frac{n-1}{2}}W_{L}^{\frac{n-1}{2}}}{V_{R}^{\frac{n-1}{2}}W_{R}^{\frac{n-1}{2}}}(\epsilon-\text{i}u)\frac{\sin(n\theta)}{\sin(\theta)},$ (IV.12)
$\ell_{2n-1}$ $=$ $\frac{V_{L}^{\frac{n-1}{2}}W_{L}^{\frac{n-1}{2}}}{V_{R}^{\frac{n-1}{2}}W_{R}^{\frac{n-1}{2}}}\frac{\left(\sin(n\theta)+\delta^{-\frac{1}{2}}\sin((n-1)\theta))\right)}{V_{R}^{-1}\sin(\theta)}.$
Note that all the eigenstates are, in general, localized at the boundaries for the open case in contrast to the closed case where they are extended Bloch waves. More specifically, the right eigenstates are localized at the left (right) boundary for $\Gamma<1$ ($\Gamma>1$) with $\Gamma=V_{R}W_{R}/V_{L}W_{L}$. Conversely, the left eigenstates are localized at the right (left) boundary for $\Gamma<1$ ($\Gamma>1$). There is no localization if $\Gamma=1$. The localization of all the eigenstates when switching from closed to open boundary conditions is called the non-Hermitian skin effect.
IV.2 Gap closing and exceptional points
Similar to the case with closed boundary conditions, one can define a complex gap
$\Delta_{c}^{(OBC)}=(\epsilon^{+}(\theta)-\epsilon^{-}(\theta))/2=\sqrt{\tilde{H}_{1}(\theta)\tilde{H}_{2}(\theta)-u^{2}}$
$=\!\!\sqrt{\!V_{L}V_{R}+W_{L}W_{R}+2\sqrt{V_{L}}\sqrt{V_{R}}\sqrt{W_{L}}\sqrt{W_{R}}\cos\theta-u^{2}}$
as well as a minimal gap
$\Delta^{(OBC)}=\text{min}_{\theta}|\Delta_{c}^{(OBC)}(\theta)|,$ (IV.15)
where $\theta$ is given by (IV.8) or by a solution of (IV.9) depending on the parity of $N$. A different analysis of the complex gap is required for each parity of $N$.
IV.2.1 Odd N
Let us start with odd $N$ where the parameter $\theta\rightarrow\theta_{m}$ is explicitly given by (IV.8). Also, for clarity, let us consider first $u=0$. In this case, the complex gap (IV.2) will vanish if $\tilde{H}_{1}(\theta_{m})\tilde{H}_{2}(\theta_{m})=0$, which happens if any of the hopping parameters vanishes or if
$\left(\sqrt{\delta}\right)_{\text{EP}}=-e^{-\text{i}\theta_{m}}~{}~{}\text{or}~{}~{}\left(\sqrt{\delta}\right)_{\text{EP}}=-e^{\text{i}\theta_{m}}.$ (IV.16)
Therefore, for non-vanishing hopping parameters, there are $N-1$ exceptional points. For finite $N$, these points always have an imaginary part. In the thermodynamic limit, however, the exceptional points will converge to the real axis
$\lim_{N\rightarrow\infty}\left(\sqrt{\delta}\right)_{\text{EP}}=\pm 1$ (IV.17)
in the cases when $\theta_{m}\to 0$ or $\theta_{m}\to\pi$.
Similarly to the closed boundary case, turning on the field $u$ couples $\tilde{H}_{1}(\theta_{m})$ and $\tilde{H}_{2}(\theta_{m})$. We observe that introducing the rescaled field,
$\bar{u}=u/(\sqrt{W_{L}}\sqrt{W_{R}})$ (IV.18)
one can write
$\tilde{H}_{1}(\theta_{m})\tilde{H}_{2}(\theta_{m})-u^{2}=\sqrt{V_{L}}\sqrt{V_{R}}\sqrt{W_{L}}\sqrt{W_{R}}$
×$\quad\times\left(\sqrt{\delta}+\frac{1}{\sqrt{\delta}}(1-\bar{u}^{2})+2\cos(\theta_{m})\right)\,.$ (IV.19)
The minimal gap (IV.15) as a function of $\bar{u}$ and $\delta$ is shown in Fig. 3.
Figure 3: The rescaled minimal gap $\Delta$ as a function of $\sqrt{\delta}$ and $\bar{u}$ for $N=101$ and open boundary conditions.
The exceptional points are therefore now given by
$\left(\sqrt{\delta}\right)_{\text{EP}}=-\cos(\theta_{m})\pm\sqrt{\bar{u}^{2}-\sin^{2}(\theta_{m})}.$ (IV.20)
Interestingly, the expression (IV.20) implies that real exceptional points do exist for finite $N$ provided that $\bar{u}\geq|\sin(\theta_{m})|$ or $\bar{u}\leq-|\sin(\theta_{m})|$. In the thermodynamic limit, critical fields $\bar{u}\rightarrow\pm 1$ emerge when $\theta_{m}$ approaches 00 or $\pi$. In Fig. 4 some examples of the loci of exceptional points for different values of the rescaled field $\bar{u}$ are shown.
Figure 4: Loci of exceptional points $\left(\sqrt{\delta}\right)_{\text{EP}}$ for $N=41$ and different rescaled fields $\bar{u}$.
Recall that when $N$ is odd, there is, in addition, an isolated root $\epsilon=\text{i}u$. We can impose that this root coincides with the quasienergies (IV.4), that is, $\text{i}u=\epsilon^{\pm}$. For non-zero hopping parameters, this implies again the condition (IV.16), but now associated with the degenerate quasienergy $\text{i}u$.
IV.2.2 Even N
Similar to the previously discussed case of odd $N$, the complex gap (IV.2) will vanish if
$\sqrt{\delta}+\frac{1}{\sqrt{\delta}}(1-\bar{u}^{2})+2\cos(\theta)=0$ (IV.21)
but with $\theta$ now given by by (IV.9). Since Eq. (IV.9) also involves $\sqrt{\delta}$, one has to consider both equations simultaneously. Note that both conditions are satisfied if any of the hopping parameters vanishes similar to the case of $N$ odd. Otherwise, after some manipulation, the intersection of (IV.21) and (IV.9) leads to,
$\sin(\theta)\pm\bar{u}\sin\left(\left(\frac{N}{2}+1\right)\theta\right)=0,$ (IV.22)
whose solutions $\theta=\theta_{EP}$ determine exceptional values of $\sqrt{\delta}$ from either (IV.21) or (IV.9). That is, choosing (IV.9), the exceptional points are given by,
$\left(\sqrt{\delta}\right)_{\text{EP}}=-\frac{\sin\left(\frac{N}{2}\theta_{\text{EP}}\right)}{\sin\left((\frac{N}{2}+1)\theta_{\text{EP}}\right)}.$ (IV.23)
The transcendental equations (IV.22) can be solved numerically for different rescaled fields $\bar{u}$. We observe that each sign in (IV.22) has in general $N/2$ solutions with $\text{Re}(\theta_{\text{EP}})\geq 0$ which determine $N$ exceptional points (other solutions with $\text{Re}(\theta_{\text{EP}})<0$ do not produce new exceptional points). An exception is the case $\bar{u}=1$ where the number of solutions of Eq. (IV.22) with a negative sign has one less solution. Nevertheless, all distinct exceptional points are obtained. We show some examples in Fig. 5.
Figure 5: In the upper panel, solutions of the transcendental equation (IV.22) for different fields $\bar{u}$ and $N=40$ are shown. In the middle panel, we fix $\bar{u}=1/2$ and consider different lattice sizes. In the bottom panel, we show the exceptional points associated with (IV.22).
There is an additional possibility to cancel the discriminant, first discovered in the study of exceptional points of the Baxter free parafermionic model [50]. It consists in imposing that the derivative of (IV.9) vanishes,
$\sqrt{\delta}(N+2)\cos\left(\left(\frac{N}{2}+1\right)\theta\right)+N\cos\left(\frac{N}{2}\theta\right)=0$ (IV.24)
in addition to (IV.9). The intersection of (IV.24) and (IV.9) then gives [50],
$\sin((N+1)\theta)-(N+1)\sin(\theta)=0$ (IV.25)
whose solutions $\theta=\theta_{\text{EP}}^{\prime}$ give extra exceptional points
$\left(\sqrt{\delta}\right)_{\text{EP}}=-\frac{\sin\left(\frac{N}{2}\theta_{\text{EP}}^{\prime}\right)}{\sin\left((\frac{N}{2}+1)\theta_{\text{EP}}^{\prime}\right)}.$ (IV.26)
The transcendental equation (IV.25) is clearly independent of the field $u$. It can be solved numerically; there are $N-2$ solutions [50] with $\text{Re}(\theta_{\text{EP}}^{\prime})\geq 0$ which fully describe the field independent part of the discriminant. The numerical solution for different lattice sizes is shown in the upper panel of Fig. 6. Note that in this case the associated repeated quasienergy is not zero. Instead, they trace interesting curves in the complex plane, see the examples shown in the bottom panel of Fig. 6.
Figure 6: (a) The solution of Eq. (IV.26) leads to additional, $\bar{u}$ independent exceptional points. (b) The corresponding energies $\varepsilon_{EP}$ lie along non-trivial curves in the complex plane.
V Bulk-boundary correspondence
It is well-known that the standard bulk-boundary correspondence for topological Hermitian systems breaks down in the non-Hermitian case. Mathematically, the fundamental issue is that the eigenvalues of a finite non-Hermitian matrix do, in general, not converge to the eigenvalues of the semi-infinite case. This is similar to a discontinuous point. More specifically, in the Hermitian case we can connect the topological invariant calculated for a periodic system with the number of edge states in a finite system. The energy of these edge states converges exponentially to zero with system size. This is what is known as the standard bulk-boundary correspondence.
In the non-Hermitian case, the eigenspectrum cannot be used to construct a bulk boundary correspondence for a finite system. Instead, it is known from Toeplitz theory that the singular value spectrum has to be used. The K-splitting theorem states [51] that a system with winding number $I$ has $K\geq|I|$ singular values which will converge to zero with increasing system size and which belong to states which are exponentially localized at the boundary. In the semi-infinite case these states become exact eigenstates with exactly zero energy but for finite system size they only get mapped exponentially close to zero by the Hamiltonian but are not exact eigenstates. I.e., in a finite non-Hermitian system the zero energy edge modes are ’hidden’, in general, and constitute very long-lived states. The proper bulk-boundary correspondence for non-Hermitian systems was recently fully explored in Ref. [25]. We note that this formalism does encompass the Hermitian case because for a Hermitian matrix the singular values are just the absolute values of the eigenvalues. Here we want to demonstrate the bulk-boundary correspondence explicitly for the model under consideration.
V.1 Singular values
To do so, we start by computing the eigenvalues of
$\Sigma_{N}=\mathcal{T}_{N}\mathcal{T}_{N}^{\dagger}$ (V.1)
when $\gamma=u=0$, that is, the open case with zero field. The singular values $\sigma$ of $\mathcal{T}_{N}$ are given by the square roots of the eigenvalues of $\Sigma_{N}$ which is a symmetric band matrix with elements,
$(\Sigma_{N})_{{2j-1,2j+1}}=(\Sigma_{N})_{{2j-1,2j-3}}=V_{L}W_{R},$
$(\Sigma_{N})_{{2j,2j-2}}=(\Sigma_{N})_{{2j,2j+2}}=W_{L}V_{R},$
$(\Sigma_{N})_{{2j-1,2j-1}}=V_{L}^{2}+W_{R}^{2},~{}~{}(\Sigma_{N})_{{2j,2j}}=W_{L}^{2}+V_{R}^{2}$
$(\Sigma_{N})_{{1,1}}=V_{L}^{2},\quad(\Sigma_{N})_{{L,L}}=\begin{cases}V_{R}^{2},&\text{even~{}L}\\ W_{R}^{2},&\text{odd~{}L}\end{cases}.$ (V.2)
This means that $\Sigma_{N}$ represents a hopping model with next-nearest neighbour interaction.
The spectral problem $(\Sigma_{N})\vec{s}=\sigma^{2}\vec{s}$ reads,
$s_{2j-3}+s_{2j+1}=\frac{\sigma^{2}-(V_{L}^{2}+W_{R}^{2})}{V_{L}W_{R}}s_{2j-1}$
$s_{2j-2}+s_{2j+2}=\frac{\sigma^{2}-(W_{L}^{2}+V_{R}^{2})}{W_{L}V_{R}}s_{2j}$ (V.3)
with boundaries fixed by the last line in (V.1). Using the method reviewed in App. A, we find that there are two solutions to the difference equations (V.1).
In the first (second) solution the even (odd) components of the vector $\vec{s}$ are zero $s_{2j}=0$ ($s_{2j-1}=0$) and the left boundary fixes the ratio $s_{3}/s_{1}$ ($s_{4}/s_{2}$). The first solution has singular values
$\sigma^{2}_{\phi}=V_{L}^{2}+W_{R}^{2}+2V_{L}W_{R}\cos(\phi)$ (V.4)
and vector components,
$\frac{s_{2j-1}}{s_{1}}=\frac{W_{R}}{V_{L}}\frac{\sin((j-1)\phi)}{\sin(\phi)}+\frac{\sin(j\phi)}{\sin(\phi)},\quad j\geq 3,$ (V.5)
$\frac{s_{3}}{s_{1}}=\frac{W_{R}}{V_{L}}+2\cos(\phi),$ (V.6)
where $\phi$ satisfies the transcendental equation,
$\sin\left(\frac{N}{2}\phi\right)\left(\frac{W_{R}}{V_{L}}+2\cos(\phi)\right)-\sin\left(\left(\frac{N}{2}-1\right)\phi\right)=0.$ (V.7)
The second solution has eigenvalues
$\sigma^{2}_{\omega}=W_{L}^{2}+V_{R}^{2}+2W_{L}V_{R}\cos(\omega)$ (V.8)
and vector components,
$\frac{s_{2j}}{s_{2}}=\frac{\sin(j\omega)}{\sin(\omega)},\quad j\geq 3,$ (V.9)
$\frac{s_{4}}{s_{2}}=2\cos(\omega)$ (V.10)
with the parameter $\omega$ satisfying,
$\sin\left(\left(\frac{N}{2}+1\right)\omega\right)+\frac{W_{L}}{V_{R}}\sin\left(\frac{N}{2}\omega\right)=0.$ (V.11)
The spectrum of (V.1) is given by the union of $\sigma_{\phi}^{2}$ and $\sigma_{\omega}^{2}$. Remarkably, the pairs of variables $\{V_{L},W_{R}\}$ and $\{V_{R},W_{L}\}$ decouple from each other. This means that the characteristic polynomial associated with (V.1) factorizes into two independent factors depending solely on each of these pairs.
The transcendental equations (V.7) and (V.11) can be solved numerically. We find that (V.7) has $N/2$ real solutions in the interval $0<\phi<\pi$ when $|V_{L}/W_{R}|\geq 1$. When $|V_{L}/W_{R}|<1$ a complex conjugate pair emerges near $\pi$. This complex root produces an exponentially small singular value $\sigma_{\phi}$. Similarly, we find that (V.11) has $N/2$ real solutions in the interval $0<\omega<\pi$ when $|V_{R}/W_{L}|\geq 1$ and $N/2-1$ real solutions plus a complex conjugate pair solution when $|V_{R}/W_{L}|<1$. In Fig. 7, we plot the smallest two singular values coming from each transcendental equation considering a lattice with $N=100$ sites.
Figure 7: The two smallest singular values from Eq. (V.4) (upper panel) and from Eq. (V.8) (lower panel). In both cases the corresponding transcendental equation is solved numerically.
The Bloch Hamiltonian for the corresponding periodic problem with $u=0$ has block structure, see Eq. (III.31). There are therefore two winding numbers $\nu_{1,2}$ defined in Eq. (III.32). The k-splitting theorem then predicts that there are $K\geq|\nu_{1}|+|\nu_{2}|$ singular values $\sigma$ with $\lim_{N\to\infty}\sigma=0$ [51, 25]. This is fully consistent with our results, compare Fig. 7 with the phase diagram in the left panel of Fig. 2.
Figure 8: The smallest two singular values (top row) and the smallest two eigenvalues (bottom row) for $W_{L}=0.5$ (left column) and $W_{L}=1.5$ (right column) with $V_{R}=W_{R}=1$ for a system with $N=200$. The two eigenvalues are always degenerate. The solid lines are for the unperturbed system, the dashed lines for a system where a random matrix with matrix elements $|a_{ij}|<0.02$ has been added. The numbers in the top row indicate the sum of the winding numbers $|\nu_{1}|+|\nu_{2}|$.
V.2 Topological protection and hidden zero modes
In the limit of a semi-infinite chain, the system will have $K\geq|\nu_{1}|+|\nu_{2}|$ topologically protected edge modes with exactly zero energy. However, these states will only be eigenstates in this limit but not for a finite system. For a finite system, the singular values have to be considered as we have done above. Here we want to show explicitly how the exact edge states in the semi-infinite limit arise and that they become ’hidden’ if the system is finite.
For a semi-infinite chain, we can find zero modes by looking for solutions of $Hv=0$ and $\tilde{H}w=0$. Here $\tilde{H}=H(h_{j}\to h_{-j})$ is the reflected Hamiltonian with Fourier components $h_{j}$. The reason that we have to consider also the reflected Hamiltonian is that we are thinking about a system which extends up to infinity either from an edge to the left or to the right.
The condition $Hv=0$ leads to $v_{2}=0$ and $W_{R}v_{2j}+V_{L}v_{2j+2}=0$ implying that all even vector components are zero, $v_{2j}=0$. For the odd vector components we find the recurrence relation $V_{R}v_{2j-1}+W_{L}v_{2j+1}=0$ with the solution $v_{2j+1}=\left(-\frac{V_{R}}{W_{L}}\right)^{j}v_{1}$. In addition, we have to demand that the solution is normalizable, $||v||^{2}=|v_{1}|^{2}\sum_{j=0}^{\infty}\left(-\frac{V_{R}}{W_{L}}\right)^{2j}=1$. This is possible if $|W_{L}|>|V_{R}|$. In this case, the semi-infinite chain has a zero energy edge mode $v$ exponentially localized at the left edge with vector components
$v_{2j+1}=\sqrt{1-\left(\frac{V_{R}}{W_{L}}\right)^{2}}\left(-\frac{V_{R}}{W_{L}}\right)^{j},\quad v_{2j}=0\,.$ (V.12)
Note that the only solution of $Hv=0$ for a finite system is the trivial vector $v=0$.
Similarly, we can consider zero modes of the reflected Hamiltonian, $\tilde{H}w=0$. In this case we find that an exponentially localized zero mode exists if $|W_{R}|>|V_{L}|$ and has vector components
$w_{2j+1}=0,\quad w_{2j+2}=\sqrt{1-\left(\frac{V_{L}}{W_{R}}\right)^{2}}\left(-\frac{V_{L}}{W_{R}}\right)^{j}\,.$ (V.13)
Again, for a finite system the only solution to $\tilde{H}w=0$ is $w=0$. The non-trivial edge mode with energy zero only exists in the thermodynamic limit. We note that we find the exact same separation of parameters as for the singular values: the first zero mode depends only on $\{V_{R},W_{L}\}$ and the second zero mode only on $\{V_{L},W_{R}\}$. For a finite system, we can truncate the vectors $v,w$. They are then no longer eigenvectors, $Hv\neq 0$, but they do get mapped exponentially close to zero, i.e. $||Hv||\to 0$ for $N\to\infty$ and similarly for $w$. This is what we mean by hidden zero modes.
For the example considered in Fig. 8 with $V_{R}=W_{R}=1$ this means that the first zero mode is stable for $-1<V_{L}<1$ and the second zero mode for $|W_{L}|>1$. This is fully consistent with the singular values as well as with the winding numbers as is expected based on the K-splitting theorem. We thus do have explicitly confirmed the bulk-boundary correspondence in this case.
To check that the singular values with $\lim_{N\to\infty}\sigma=0$ are topologically protected while zero eigenvalues are, in general, not protected for a finite non-Hermitian system, we also show in Fig. 8 both quantities when calculated for a finite Hamiltonian $H$ ($N=200$) with a random complex matrix $A$ with elements $|a_{ij}|\leq 0.02$ added. The results show that the singular values are stable and thus indeed topologically protected while the eigenvalues are, in general, not protected. More specifically, we find that for the case $W_{L}=0.5$ shown in the lower left panel there are two degenerate zero eigenvalues in the unperturbed system for $-0.4\lesssim V_{L}\lesssim 0.5$. However, the small perturbation immediately moves them away from zero energy. They are not protected.
The situation is slightly more complicated in the case $W_{L}=1.5$ shown in the lower right panel of Fig. 8. In the unperturbed case, there are two degenerate zero energy eigenvalues for $-1.3\lesssim V_{L}\lesssim 1.5$. Interestingly, adding a perturbation does not completely remove them from zero energy. Instead, they are stable for $-1<V_{L}<1$. This can be understood as follows: in this regime, the non-Hermitian Hamiltonian with $V_{R}=W_{R}=1$ and $W_{L}=1.5$ can be adiabatically connected to a chiral Hermitian Hamiltonian without closing the gap. This can be achieved, for example, by $W_{R}\to 1.5$ and then $V_{R}\to V_{L}$. For the obtained Hermitian Hamiltonian one can calculate the winding number of the upper block which turns out to be $I=1$. The standard bulk-boundary correspondence then predicts the existence of two protected zero-energy edge modes. Since the gap never closes and the sub-lattice symmetry remains intact, these zero modes survive in the non-Hermitian case. However, the full topology of the non-Hermitian model also in this case is only captured by the singular values, not by the eigenvalues.
VI Entanglement entropy
In this section we investigate the entanglement entropy for the non-Hermitian model (II) with periodic boundary conditions. We use the reduced density matrix approach for free fermions [52, 53] which has been extended to the non-Hermitian realm [27, 28, 54, 26, 55] using the generalized density matrix $\rho=|r\rangle\langle\ell|$ where $\langle\ell|$ and $|r\rangle$ are bi-orthogonal left and right eigenstates of (II). The key quantity is the two-point correlation matrix with elements
$C_{ij}=\langle\ell|c_{i}^{\dagger}c_{j}|r\rangle={\rm tr}\,\left(\rho c_{i}^{\dagger}c_{j}\right)$ (VI.1)
Restricting the indices to a given subsystem $\mathcal{A}$, the von Neumann entanglement entropy is given by,
$S_{\mathcal{A}}=-\sum_{j}\nu_{j}\log\nu_{j}+(1-\nu_{j})\log(1-\nu_{j})$ (VI.2)
where $\nu_{j}$ are the eigenvalues of the correlation matrix (VI.1) restricted to $\mathcal{A}$. In this paper we only consider $\mathcal{A}=\{1,\dots,N/2\}$, that is, we cut the system in half. For the model considered in this paper, in contrast to Hermitian systems, the eigenvalues $\nu_{j}$ and $S_{\mathcal{A}}$ can be complex. Nevertheless, we will indiscriminately call $S_{\mathcal{A}}$ the entanglement entropy. In addition, we choose the branch cut of the logarithm to be on the negative real axis.
Also, in non-Hermitian systems, the notion of a ground state is not well defined since the spectrum is complex in general. We have to distinguish two cases. If there is a line gap, then we fill the states on one side of the line gap so that there is an excitation gap. If there is no line gap then constructing a ground state is somewhat arbitrary but we will describe in each case what the chosen ground state is.
To evaluate (VI.1), we first organize the eigenvalues in the desired ordering,
$D=\text{diag}(\epsilon(k_{\sigma(1)}),\dots,\epsilon(k_{\sigma(N)}))$ (VI.3)
where $\epsilon(k_{\sigma(j)})$ denotes one of the $N$ possible quasienergies. Accordingly, define the column vectors,
$\vec{r}(k_{\sigma(j)})=\sqrt{\frac{2}{N}}\begin{pmatrix}r_{1}(k_{\sigma(j)})\\ r_{2}(k_{\sigma(j)})\\ \vdots\\ r_{N}(k_{\sigma(j)})\\ \end{pmatrix},$ (VI.4)
$\vec{\ell}(k_{\sigma(j)})=\sqrt{\frac{2}{N}}\begin{pmatrix}\ell_{1}(k_{\sigma(j)})\\ \ell_{2}(k_{\sigma(j)})\\ \vdots\\ \ell_{N}(k_{\sigma(j)})\\ \end{pmatrix},$ (VI.5)
with components given by (III.1,III.2), which are respectively right and left eigenvectors of $\mathcal{T}_{N}$ with eigenvalue $\epsilon(k_{\sigma(j)})$. Introduce the matrices
$R=\left(\vec{r}(k_{\sigma(1)}),\dots,\vec{r}(k_{\sigma(N)})\right),$ (VI.6)
$L=\left(\vec{\ell}(k_{\sigma(1)}),\dots,\vec{\ell}(k_{\sigma(N)})\right).$ (VI.7)
It follows that,
$\mathcal{T}_{N}R=RD,\quad L^{T}\mathcal{T}_{N}=DL^{T},\quad RL^{T}=L^{T}R=\mathds{1}.$ (VI.8)
The diagonalized hopping matrix is then given by
$\mathcal{T}_{N}=\bm{\eta}^{\dagger}D\bm{\eta},\quad\bm{\eta}=L^{T}\bm{c},\quad\bm{\eta}^{\dagger}=\bm{c}^{\dagger}R$ (VI.9)
implying the algebra (II.2) also for the $\bm{\eta}$. We have,
$c_{j}=\sum_{m=1}^{N}R_{jm}\eta_{m},\quad c_{i}^{\dagger}=\sum_{n=1}^{N}L_{in}\eta_{n}^{\dagger}.$ (VI.10)
Finally, define the states,
$\langle\ell|=\langle 0|\prod_{a\in occ.}\eta_{a},\quad|r\rangle=\prod_{a\in occ.}\eta_{a}^{\dagger}|0\rangle,$ (VI.11)
and compute,
$C_{ij}=\langle\ell|c_{i}^{\dagger}c_{j}|r\rangle=\sum_{a\in occ.}L_{ia}R_{ja}.$ (VI.12)
We now investigate the spectrum of the correlation matrix (VI.12) and the associated entanglement entropy for indices in (VI.12) restricted to $i,j\in\{1,\dots,N/2\}$. We perform numerical calculations for large lattices, in which case we only keep eigenvalues of the correlation matrix that satisfy $|\nu_{j}|>10^{-50}$ and $|1-\nu_{j}|>10^{-50}$. Additionally, we set to zero real or imaginary parts of $\nu_{j}$ and $1-\nu_{j}$ that are individually smaller than $10^{-50}$. This is particular important in cases where the eigenvalues of the correlation matrix are negative with a tiny imaginary part, whose sign can numerically oscillate around the branch cut. Before considering numerics, as a warm-up, we study the trivial 2-cell model ($N=4$) analytically. Both zero field $u=0$ and $\mathcal{PT}$-symmetric cases are considered.
VI.1 Sub-lattice symmetric case
For the zero field case, setting $V_{R}=W_{R}=1$, let us consider transitions between the phases $(\nu_{1},\nu_{2})=(1,-1)$ (with $W_{L}>1$) and $(0,-1)$, between $(1,0)$ and $(0,0)$ and between $(1,-1)$ (with $W_{L}<-1$) and $(0,-1)$, recall the phase diagram shown in the left panel of Fig. 2. For concreteness, let us consider only $V_{L}\geq 0$. Recall that the spectrum of $\mathcal{T}_{N}$ in the phases $(1,-1)$ and $(0,0)$ is in general composed of two lobes in the complex plane either separated by the imaginary axis or real axis, therefore being characterized by $I_{1}=0$. On the other hand, the spectrum in the phases $(1,0)$ and $(0,-1)$ traces a closed path around $E_{B}=0$, therefore being characterized by $I_{1}\neq 0$.
We start with the 2-cell model ($N=4$). For this lattice size, there is no lobe in the spectrum, and the eigenvalues are either purely real or purely imaginary. We fill the states with energy $-\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$ and leave empty the states with energy $\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$. The associated correlation matrix is given by,
$C=\left(\begin{array}[]{cc}\frac{1}{2}&-\frac{\sqrt{\left(V_{L}-1\right)\left(1-W_{L}\right)}}{4\left(V_{L}-1\right)}-\frac{\sqrt{\left(V_{L}+1\right)\left(W_{L}+1\right)}}{4\left(V_{L}+1\right)}\\ -\frac{\sqrt{\left(V_{L}-1\right)\left(1-W_{L}\right)}}{4\left(1-W_{L}\right)}-\frac{\sqrt{\left(V_{L}+1\right)\left(W_{L}+1\right)}}{4\left(W_{L}+1\right)}&\frac{1}{2}\\ \end{array}\right),$ (VI.13)
which is non-Hermitian and can have complex eigenvalues. Indeed, we plot both real and imaginary parts of the two eigenvalues of (VI.13) in Fig. 9 for some fixed values of $W_{L}$ as function of $V_{L}$.
We observe that within the phases characterized by $(1,-1)$ the eigenvalues form conjugate pairs of the form $1/2\pm\text{i}\alpha$, see panels (a) and (c) in Fig. 9 for the fixed values $W_{L}=\pm 1.5$. The same form of eigenvalues is observed for other values of $|W_{L}|>1$. Thus, in this phase, the entanglement entropy is real and given by,
$S_{2}=-\log\left(\frac{1}{4}+\alpha^{2}\right)+4\alpha\arctan(2\alpha)$ (VI.14)
where $\alpha$ depends on $V_{L}$ and $W_{L}$. We also note that $\alpha\rightarrow\infty$ as the transition point $V_{L}=1$ is approached and, as a consequence, $S_{2}$ diverges at $V_{L}=1$. On the other hand, deep inside the phase, taking $V_{L}=0$ and $W_{L}\rightarrow\infty$, we have $\alpha\rightarrow 0$ and therefore $S_{2}\rightarrow 2\log 2$.
Increasing $V_{L}$ and crossing the transition point $V_{L}=1$, we observe that the eigenvalues of the correlation matrix in the phase $(0,-1)$ are complex, with the form $\beta_{1,2}\pm\text{i}\alpha$. That is, the eigenvalues have imaginary parts with opposite signs but with different real parts $\beta_{2}=1-\beta_{1}$. This leads to complex values for the entanglement entropy. However, for large $|W_{L}|$ the imaginary part tends to disappear when $V_{L}$ increases. In the limit $V_{L}\rightarrow\infty$ and $|W_{L}|\rightarrow\infty$, the exact eigenvalues of the correlation matrix are $41/2\pm\sqrt{2}/$ 4, producing a real entropy $S_{2}\approx 0.832991$.
In the phase $(1,0)$, the eigenvalues once more have the form $\beta_{1,2}\pm\text{i}\alpha$ with $\beta_{2}=1-\beta_{1}$, and the entanglement entropy is complex (except if $V_{L}=W_{L}$), see panel (b) in Fig. 9 for the fixed value $W_{L}=0.1$. Moving to the phase $(0,0)$, we observe that the eigenvalues of $C$ are real, however, they are greater than unity or smaller then zero, and therefore also produce a complex entanglement entropy. Only in the limit $V_{L}\rightarrow\infty$ and $W_{L}\rightarrow 0$, do the eigenvalues of the correlation matrix tend to $1$ and 00, thus producing a zero entanglement entropy.
Figure 9: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the trivial two-cell model ($N=4$) with zero field $u=0$, $W_{R}=V_{R}=1$ and some fixed values of $W_{L}$ indicated in the labels (a)-(c) as a function of $V_{L}$.
Now, we consider larger lattice sizes using numerical diagonalization. We observe a variety of possible types of eigenvalues in the correlation spectrum, similar to the trivial 2-cell model. In some phases, we may find eigenvalues with $\text{Re}(\nu_{j})>1$ as well as $\text{Im}(\nu_{j})\neq 0$. Also, we note that near the transition points the behaviour of the correlation matrix eigenvalues is more intricate. Let us analyze each phase in details. Results are collected in Fig. 10 for a lattice composed of $N=120$ sites. In the insets of Fig. 10, we show the typical form of the complex spectrum indicating the filling adopted, as well as the real part of the eigenvalues of the correlation matrix, considering a smaller lattice with $N=52$ for better visualization of the results.
Figure 10: Entanglement entropy for the model with sublattice symmetry (zero field $u=0$) for $V_{R}=W_{R}=1$ and some fixed values of $W_{L}$ indicated in the labels (a)–(c), as a function of $V_{L}$. The lattice size is $N=120$. In the insets, the typical form of the complex spectrum is shown; the occupied states are indicated by filled circles while the unoccupied states by open circles. The typical eigenvalues of the correlation matrix are represented by empty up triangles (only the real part is shown since the complex part when present is small). For better visualization, in the insets we consider $N=52$.
[…]
Moving to the phases $(0,-1)$, see once more panels (a) and (c) in Fig. 10, we note that all eigenvalues of the correlation matrix satisfy $10<\text{Re}(\nu_{j})<$ 1, with most eigenvalues being either close to unity or zero. Some intermediate eigenvalues appear, and these carry a small imaginary part. They produce a finite entanglement entropy, also complex, but with a small imaginary part. However, the number of intermediate states actually grows with the length of the chain. This implies a scaling of the entanglement entropy, see Fig. 11, where we fix $W_{L}=1.5$ and some values of $V_{L}$ deep inside the phase. We observe a scaling of the form $\text{Re}(S_{N})\sim(c/3)\log(N)$ with $c=1$, as well as a decay $\text{Im}(S_{N})\sim\exp(-\log(N))$ of the imaginary part. It follows that, remarkably, this phase may be described by a CFT with central charge $c=1$, even though it has a point gapped spectrum. Similar results can be obtained for $W_{L}<-1$ in the same phase $(0,-1)$.
Figure 11: Scaling of the entanglement entropy in the phase $(0,-1)$ for $W_{L}=1.5$ and values of $V_{L}$ deep inside the phase.
The phase $(1,0)$, see panel (b) in Fig. 10, is similar to the phases $(0,-1)$. The scaling of the entanglement entropy for this case is shown in Fig. 12. Also in this phase the scaling of the entanglement entropy is consistent with a central charge $c=1$.
Figure 12: Scaling of the entanglement entropy in the phase $(1,0)$ for $W_{L}=0.1$ and some values of $V_{L}$ inside the phase. Note that for $V_{L}=0$ the imaginary part is negative, so it is its absolute value that decreases with the length of the chain.
The last phase to be discussed in this subsection is the phase $(0,0)$. Here, we observe eigenvalues of the correlation matrix with $\text{Re}(\nu_{j})>1$ and also some eigenvalues with a small imaginary part. We note that, for fixed parameters $W_{L}$ and $V_{L}$ in this phase, the imaginary part of the entanglement entropy decreases with the length of the chain, while the real part slightly increases and saturates. However, as we go deeper into this phase by increasing $V_{L}$, the entanglement entropy tends to disappear. This phase is therefore trivial.
In summary, for the zero field case and in the thermodynamic limit, the entanglement entropy is real deep inside the quantum phases, and discontinuous near the transition points. The gapped non-trivial phases have two protected eigenvalues at $1/2$, contributing $S_{N}=2\log 2$ to the entanglement entropy.
VI.2 $\mathcal{PT}$-symmetric case
Let us now consider the $\mathcal{PT}$ symmetric case with PBC, setting $u=1$. For simplicity, we will restrict ourselves to the quadrant with $V\geq 0$ and $W\geq 0$ in the phase diagram shown in the right panel of Fig. 2. Then, four regions must be analyzed. The region $|W-V|<1<W+V$ is the exceptional phase, characterized by a large number of exceptional points (recall Sec. III.2). In this phase, the $\mathcal{PT}$-symmetry is broken. The regions $W-V<-1$ and $W-V>1$ are adiabatically connected to the standard SSH model via $u\rightarrow 0$. Then, the former is called the $\mathcal{PT}$-unbroken trivial phase and the latter the $\mathcal{PT}$-unbroken topological phase. In the phase $W+V<1$ the spectrum is purely imaginary, and we name it the complex phase.
As before, we start our analysis by the simple 2-cell model ($N=4$), which has the following correlation matrix,
$C=\frac{1}{4}\left(\begin{array}[]{cc}\frac{(V-W)^{2}}{(V-W)^{2}-1+\text{i}\sqrt{(V-W)^{2}-1}}+\frac{(V+W)^{2}}{(V+W)^{2}-1+\text{i}\sqrt{(V+W)^{2}-1}}&\frac{W-V}{\sqrt{(V-W)^{2}-1}}-\frac{V+W}{\sqrt{(V+W)^{2}-1}}\\ \frac{W-V}{\sqrt{(V-W)^{2}-1}}-\frac{V+W}{\sqrt{(V+W)^{2}-1}}&\frac{(V-W)^{2}}{(V-W)^{2}-i\sqrt{(V-W)^{2}-1}-1}+\frac{(V+W)^{2}}{(V+W)^{2}-i\sqrt{(V+W)^{2}-1}-1}\\ \end{array}\right)$
associated with the filled states $-\sqrt{(V\mp W)^{2}-1}$ and empty states $\sqrt{(V\mp W)^{2}-1}$. We plot in Fig. 13 the eigenvalues of $C$ and the associated entanglement entropy for some fixed values of $V$ as a function of $W$, covering all the four regions of the phase diagram.
For $V=0$ and $10<W<$ 1 that is, within the complex phase, see panel (a) in Fig. 13, we can observe that the eigenvalues of the correlation matrix are real with $\nu_{j}>1$ or $\nu_{j}<0$ until the phase transition point $W=1$ is reached. As a consequence, the entanglement entropy is complex in this phase. Crossing the transition point ($W=1$) towards the $\mathcal{PT}$-unbroken topological phase, we note that the eigenvalues have the form $1/2\pm\text{i}\alpha$, and the entanglement entropy is given by (VI.14). In the limit $W\rightarrow\infty$, $\alpha$ becomes negligible and the entanglement entropy tends to $S_{2}\rightarrow 2\log(2)$, indeed characteristic of a topological phase with winding $I=1$.
Increasing $V$ to $V=0.5$, see panel (b) in Fig. 13, the behaviour of the eigenvalues $\nu_{j}$ is the same as for $V=0$ until the exceptional phase is reached at $W=0.5$. Within the exceptional phase, the eigenvalues $\nu_{j}$ have the form $\beta_{1,2}\pm\text{i}\alpha$, that is, the eigenvalues have the same imaginary part with opposite signs but different real parts related by $\beta_{2}=1-\beta_{1}$. As a consequence, the entanglement entropy is complex in the exceptional phase. This phase ends at $W=1.5$ from which point the $\mathcal{PT}$-unbroken topological phase is again reached being characterized by a real entanglement entropy which tends to $S_{2}\rightarrow 2\log(2)$ deep inside the phase.
To conclude the analysis of the 2-cell model, we consider $V=1.5$ such that the $\mathcal{PT}$-unbroken trivial phase is considered. This phase for $V=1.5$ occurs when $0.50<W<$ 0.5, and we note that there the eigenvalues are real with $\nu_{j}>1$ or $\nu_{j}<0$, producing again a complex entanglement entropy except if $W=0$. However, deep inside the phase ($V\rightarrow\infty$), the eigenvalues tend to 00 and $1$, leading to a null entanglement, as expected from a trivial phase. For $2.50.5<W<$ 2.5 we encounter once more the exceptional phase and for $W>2.5$ the $\mathcal{PT}$-unbroken topological phase.
Figure 13: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the trivial $\mathcal{PT}$-symmetric two-cell model ($N=4$) with $u=1$, some fixed values of $V$ indicated in the labels (a)-(c) as a function of $W$.
Finally, we consider the $\mathcal{PT}$ case for large lattice sizes. Results for the correlation spectrum are collected in Fig. 14 (where $N=52$) and for the entanglement entropy in Fig. 15 (where $N=120$). The same four regions considered for the 2-cell model are analyzed.
First, let us consider $V=0$ and vary $W$. For $10<W<$ 1, we note that the correlation matrix has real eigenvalues, most of them 00 or $1$, but there is one eigenvalue $\nu_{\text{max}}>1$ and one eigenvalue $\nu_{\text{min}}<0$. Remarkably, these eigenvalues do not depend on the lattice size, and therefore correspond to the two eigenvalues of the 2-cell model. For instance, for $V=0,W=0.5$, the eigenvalues are $\nu_{\text{max}}\approx 1.077350$ and $\nu_{\text{min}}\approx-0.077350$, which produces a finite entanglement entropy with negative real part and positive imaginary part, precisely the same as the 2-cell model, recall panel (a) in Fig. 13. Crossing the transition point $W=1$, we note that most of the eigenvalues of the correlation matrix are 00 or $1$, but there are a couple of eigenvalues of the form $1/2\pm\text{i}\alpha$, where $\alpha$ also does not depend on the lattice size. However, we note that for a fixed length $\alpha\rightarrow 0$ when $W\rightarrow\infty$. Therefore, like in some of the previously considered cases, the entanglement approaches $S_{N/2}\rightarrow 2\log(2)$ deep inside this phase.
Now we consider $V\neq 0$. Similarly to the 2-cell model, by fixing $10<V<$ 1, we can reach the complex phase, the exceptional phase and the $\mathcal{PT}$-unbroken topological phase by varying $W$. On the other hand, by fixing $V>1$ and varying $W$ we can reach the $\mathcal{PT}$-unbroken trivial phase, the exceptional phase and the $\mathcal{PT}$-unbroken topological phase.
Let us consider once more $V=0.5$. Within the complex phase ($0.50<W<$ 0.5 for $V=0.5$), all the eigenvalues of the correlation matrix jump above unity or below zero, see panel (a) in Fig. 14 where the parameters $W=0.45$ and $V=0.5$ are used. This results in a complex value for $S_{N/2}$, see panel (a) in Fig. 15. We note that two eigenvalues stand out, being significantly greater than $1$ or less than 00. All the eigenvalues of the correlation matrix in this phase seem to vary slowly with the size of the chain, actually tending to stabilize. For example, the difference between the largest eigenvalues for the lattices with $N=52$ and with $N=120$ is of the order of $10^{-9}$. Therefore, the entanglement entropy in this phase seems to be finite. However, it is difficult to verify this property since the model parameters in this phase lie in a finite interval. Also, the phase is surrounded by critical lines that drive the entanglement entropy to high values.
Moving to the $\mathcal{PT}$-unbroken topological phase, we note that the correlation spectrum is composed by eigenvalues close to $1$ and 00, with some of them being slightly bigger than $1$ or smaller then 00, as well as a conjugate pair of the form $1/2\pm\text{i}\alpha$, see panel (b) in Fig. 14 where the parameters are set to $W=1.7,V=0.5$. As the lattice size increases, we note that the eigenvalues tend to $1$ or 00, while $\alpha$ decreases slowly and seems to stabilize with the length of the chain. The associated entanglement entropy is real and finite. Going deep into the phase by increasing $W$, as before, $\alpha$ decreases and the entanglement tends to $S_{N/2}\rightarrow 2\log{2}$, see both panels (a) and (b) in Fig. 15.
In the $\mathcal{PT}$-unbroken trivial phase, the correlation spectrum is also composed mostly of 00’s and $1$’s, some of them slightly bigger than $1$ or smaller than 00. Also, some eigenvalues carry an imaginary part in the form of conjugate pairs, see panel (d) in Fig. 14 for the model parameters $W=0.5$ and $V=1.8$. Interestingly, similarly to previous cases, both real and imaginary part of the eigenvalues vary slowly with the length of the chain. Nevertheless, deep inside this phase (large $V$), the eigenvalues tend to 00 or $1$ and produces zero entanglement, as expected from a trivial phase; see panel (b) in Fig. 15.
Let us briefly discuss the exceptional phase. In this phase we observe an intricate behaviour of the correlation spectrum, characterized by various complex eigenvalues, see panel (c) in Fig. 14 where the parameters are set to $V=0.5,\,W=1.2$. This behaviour is reflected in an intricate evolution of the entanglement entropy, see panels (a) and (b) in Fig. 15. Interestingly, the entanglement entropy detects the presence of the real exceptional points (indicated by dashed lines in Fig. 15). For clarity, we only pick a few values of $W$ between the exceptional points. Note that in the thermodynamic limit the entanglement entropy is ill defined in the exceptional phase, since it is filled with exceptional points. For finite systems, we expect a scaling of $S_{N/2}$. This is numerically challenging and goes beyond the scope of this paper.
In brief, the entanglement entropy has a rich behaviour in the $\mathcal{PT}$ symmetric model and is capable of detecting exceptional points and phase transitions. In the $\mathcal{PT}$-unbroken topological phase there are two protected eigenvalues at $1/2$, contributing $2\log 2$ to the entanglement entropy.
Figure 14: Eigenvalues of the correlation matrix for the $\mathcal{PT}$ symmetric case with PBC and $u=1$ for some fixed values of $W$ and $V$ indicated in the labels (a)-(d). The lattice size is $N=52$. In the insets, the complex spectrum is shown; occupied states are indicated by orange filled circles and unoccupied states by blue open circles.
Figure 15: Entanglement entropy for the $\mathcal{PT}$ symmetric case with PBC and $u=1$ as a function of $W_{L}=W_{R}=W$ for two fixed values of $V_{L}=V_{R}=V$ as indicated in the labels. The lattice size is $N=120$. The dashed grid lines indicate the position of real exceptional points. In the insets, we zoom into regions where the exceptional points are very close to each other.
VII Conclusion
In this paper we have analytically studied a paradigmatic one-dimensional, non-Hermitian lattice model. Our goal was to illustrate many of the unique features of non-Hermitian systems in a concrete model and to do so largely based on analytical instead of purely numerical results. We derived the eigensystems both for closed (periodic and anti-periodic) as well as open boundary conditions. Our analytical results show how the Bloch waves in the closed case change into localized eigenstates for open boundaries thus explicitly demonstrating the non-Hermitian skin effect. Another common feature of non-Hermitian systems are exceptional points where both eigenvalues and -vectors coalesce. In the considered model, we have shown that such points typically occur when the complex gap closes and that the location of exceptional points, in general, differs between the closed and the open case. Surprisingly, we found in the $\mathcal{PT}$-symmetric phase that exceptional points can even become dense in the thermodynamic limit, forming an entire “exceptional phase”.
One of the central results of this work is an explicit demonstration of a proper bulk-boundary correspondence for a non-Hermitian system. This point has lead to some confusion in the literature. Here we have shown that the winding numbers for the periodic system determine the number of protected singular values which go to zero in the thermodynamic limit. This number also corresponds to the number of exact zero-energy edge modes for a semi-infinite chain. However, the eigenspectrum of the finite chain does, in general, not converge to that of the semi-infinite one. Thus, no bulk-boundary correspondence between winding numbers and the eigenspectrum of a finite system exists, in contrast to Hermitian systems. The exact zero-energy edge states in the semi-infinite non-Hermitian case are not exact eigenstates for a finite system. They become hidden which means that they correspond to metastable states which are, however, very long-lived and thus highly relevant for our understanding of non-Hermitian systems; see also Ref. [25] where this issue was discussed more generally. It is also important to note that even if zero-energy eigenstates for a finite system do exist, they are typically not topologically protected. For the studied model we demonstrated this explicitly by adding a small perturbation to the Hamiltonian and showing that the eigenenergies shift while the topologically protected singular values are stable.
Another aspect of topological protection in non-Hermitian systems which we explored here is the entanglement spectrum. For cases with a line gap we could, in particular, define a half-filled system where occupied and unoccupied bands are separated by a gap. Since the studied system is Gaussian, the entanglement spectrum is fully determined by the spectrum of the single-particle correlation matrix. We found that phase transitions and exceptional points are indicated by discontinuities in the entanglement entropy. Furthermore, we demonstrated that in topologically non-trivial phases there are protected eigenvalues in the entanglement spectrum which lead to lower, non-trivial entanglement bounds.
Non-trivial topology in the studied non-Hermitian system thus leads to two observable and experimentally relevant phenomena: (1) Extremely long-lived stable zero energy states, and (2) non-trivial entanglement in systems with a gap between occupied and unoccupied energy bands.
Appendix A Some details for the open boundary case
To make this paper self-contained, we review some details of the diagonalization of $\mathcal{T}_{N}$ (II) for open boundary conditions, that is, $\gamma=0$, following [44, 45].
The approach in [44] is based on recurrence relations satisfied by the characteristic polynomial,
$P_{N}(z)=\det_{N}\left(\mathcal{T}_{N}-z\right).$ (A.1)
The roots $P_{N}(z)=0$ are the quasienergies of the Hamiltonian (II.1). The characteristic polynomial (A.1) for $\gamma=0$ is special as it satisfies the recurrence relations
$P_{2j}(z)$ $=$ $(-\text{i}u-z)P_{2j-1}(z)-V_{L}V_{R}P_{2j-2}(z),$
$P_{2j-1}(z)$ $=$ $(\text{i}u-z)P_{2j-2}(z)-W_{L}W_{R}P_{2j-3}(z)$ (A.2)
depending on the parity of $N$. Note that one has an additional recurrence relation for the characteristic polynomial, namely
$P_{j}(z)$ $=$ $\left((z+\text{i}u)(z-\text{i}u)-(V_{L}V_{R}+W_{L}W_{R})\right)P_{j-2}(z)$ (A.3)
-- $V_{L}V_{R}W_{L}W_{R}P_{j-4}(z),$
where appropriate initial conditions must be chosen.
On the other hand, following [45], one notes that equations (II.6,II.7) can be iterated to the same fourth order difference equation,
$r_{j+2}+\frac{V_{R}W_{R}}{V_{L}W_{L}}r_{j-2}$
$=\left(-\frac{(\text{i}u+\epsilon)(\text{i}u-\epsilon)+(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}}\right)r_{j}.$ (A.4)
The difference equation (A) can be compared to (A.3) leading to the solution (IV.10).
In addition, following [45], introduce
$T_{n}(\alpha,\beta)=\frac{\alpha^{n+1}-\beta^{n+1}}{\alpha-\beta}$ (A.5)
and the sequence
$a_{n+2}-(\alpha+\beta)a_{n+1}+\alpha\beta a_{n}=0.$ (A.6)
Then, it follows that
$a_{n}=T_{n-2}(\alpha,\beta)a_{2}-\alpha\beta T_{n-3}(\alpha,\beta)a_{1}$ (A.7)
is the solution of the difference equation (A.6). Now choosing
$\alpha+\beta=\frac{(\text{i}u-\epsilon)(-\text{i}u-\epsilon)-(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}},$ (A.8)
$\alpha\beta=\frac{V_{R}W_{R}}{V_{L}W_{L}},$ (A.9)
it follows that
$r_{2j}=T_{j-2}(\alpha,\beta)r_{4}-\alpha\beta T_{j-3}(\alpha,\beta)r_{2},$ (A.10)
$r_{2j-1}=T_{j-2}(\alpha,\beta)r_{3}-\alpha\beta T_{j-3}(\alpha,\beta)r_{1},$ (A.11)
is a solution of (II.6,II.7,A). The boundary conditions determine,
$r_{1}$ $=$ $V_{L},\quad r_{2}=\epsilon-\text{i}u,\quad r_{3}=W_{R}+V_{L}(\alpha+\beta),$
$r_{4}$ $=$ $(\epsilon-\text{i}u)\left(2\frac{W_{R}}{V_{L}}+\alpha+\beta\right).$ (A.12)
where $r_{1}=V_{L}$ is an arbitrary choice. The conditions (A.8) give $\{\epsilon,\beta\}$ in terms of $\alpha$,
$\epsilon$ $=$ ±$\pm\sqrt{V_{L}V_{R}+W_{L}W_{R}+V_{L}W_{L}(\alpha+\beta)-u^{2}},$
$\beta$ $=$ $\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}.$ (A.13)
Using the parametrization
$\alpha=\frac{\sqrt{V_{R}}\sqrt{W_{R}}}{\sqrt{V_{L}}\sqrt{W_{L}}}\exp(\text{i}\theta),$ (A.14)
and then
$T_{n}\left(\alpha,\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}\right)=\frac{V_{R}^{\frac{n}{2}}W_{R}^{\frac{n}{2}}}{V_{L}^{\frac{n}{2}}W_{L}^{\frac{n}{2}}}\frac{\sin((n+1)\theta)}{\sin(\theta)}$ (A.15)
in (A.10,A.11) we find (IV.10,LABEL:vLodd) with the initial conditions (A).
The boundary condition at the end of the chain gives a condition on the parameter $\theta$, which depends on whether $N$ is odd or even. For odd $N$, the condition $x_{L+1}=0$ comes from (IV.10) with $12n=L+$ 1 and gives the quantized solution (IV.8). For even $N$, the condition $x_{L+1}=0$ comes from (LABEL:vLodd) with $12n-1=L+$ 1 and leads to the transcendental equation (IV.9). The solution (IV.12,LABEL:vLoddL) is obtained analogously.
Appendix B Discriminant of the characteristic polynomial
We recall that a polynomial has repeated roots if and only if its discriminant
$\text{disc}(P_{N}(z))=\text{disc}\left(\det_{N}\left(\mathcal{T}_{N}-z\right)\right)$ (B.1)
vanishes. The discriminant is given by the determinant of the Sylvester matrix [56], and it is in general an intricate expression of the parameters of the model. The vanishing of the discriminant (B.1) is a necessary condition for the existence of an exceptional point. Of course, the discriminant vanishes when the quasienergies collide. Additionally, for the model considered in this paper, the eigenvectors always coalesce when the quasienergies collide, recall the expressions (III.13,III.17,IV.10). Thus, the collision of quasienergies is also sufficient for the existence of an exceptional point.
In the main text, the conditions for gap closing/exceptional points are determined from the simple characterization of the eigenvalues by (III.7) and (IV.4). Here these results are verified by explicitly computing the discriminant (B.1) using Mathematica for different boundary conditions.
B.1 Closed boundary conditions
Let us consider $\gamma=1$ and $u=0$. For a lattice with $N=10$ sites the discriminant (B.1) is given by,
$\text{disc}(P_{10}^{\text{(PBC)}}(z))$ $=$ $10^{10}\left(W_{L}^{5}+V_{R}^{5}\right)\left(V_{L}^{5}+W_{R}^{5}\right)$ (B.2)
$\times$ $\left(V_{L}^{5}W_{L}^{5}-V_{R}^{5}W_{R}^{5}\right)^{8}.$
The roots of the factors $\left(W_{L}^{5}+V_{R}^{5}\right)\left(V_{L}^{5}+W_{R}^{5}\right)$ are clearly given by (III.26). The extra factor comes from (III.29), which can be seen by writing (III.29) like,
$\frac{V_{L}W_{L}}{V_{R}W_{R}}=e^{\text{i}(k+k^{\prime})}=e^{\frac{2}{5}\text{i}\pi(\ell+\ell^{\prime})}$ (B.3)
with $\ell+\ell^{\prime}=1,\dots,5$. As mentioned in the main text, this last factor produce degenerate eigenvalues of $\mathcal{T}_{N}$ but associated with different eigenvectors, thus they are not exceptional points.
Now let us consider again $N=10$, $\gamma=1$, but with $u\neq 0$. In this case the discriminant is given by,
$\text{disc}(P_{10}^{\text{(PBC)}}(z))$ $=$ $10^{10}\left(V_{L}^{5}W_{L}^{5}-V_{R}^{5}W_{R}^{5}\right)^{8}$
$\times$ $\left(\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}\right)U(u)$
where
$U(u)$ $=$ $u^{8}+u^{6}\left(V_{L}W_{L}+V_{R}W_{R}-4V_{L}V_{R}-4W_{L}W_{R}\right)$
$+$ $u^{4}\big{(}6V_{L}^{2}V_{R}^{2}-3V_{L}^{2}V_{R}W_{L}+V_{L}^{2}W_{L}^{2}-3V_{L}V_{R}^{2}W_{R}$
$+9V_{L}V_{R}W_{L}W_{R}-3V_{L}W_{L}^{2}W_{R}-3V_{L}W_{L}^{2}W_{R}$
$+V_{R}^{2}W_{R}^{2}-3V_{R}W_{L}W_{R}^{2}+6W_{L}^{2}W_{R}^{2}\big{)}$
$+$ $u^{2}\big{(}-4V_{L}^{3}V_{R}^{3}+3V_{L}^{3}V_{R}^{2}W_{L}-2V_{L}^{3}V_{R}W_{L}^{2}+V_{L}^{3}W_{L}^{3}$
$+3V_{L}^{2}V_{R}^{3}W_{R}-6V_{L}^{2}V_{R}^{2}W_{L}W_{R}+4V_{L}^{2}V_{R}W_{L}^{2}W_{R}$
$-2V_{L}^{2}W_{L}^{3}W_{R}-2V_{L}V_{R}^{3}W_{R}^{2}+4V_{L}V_{R}^{2}W_{L}W_{R}^{2}$
$-6V_{L}V_{R}W_{L}^{2}W_{R}^{2}+3V_{L}W_{L}^{3}W_{R}^{2}+V_{R}^{3}W_{R}^{3}$
$-2V_{R}^{2}W_{L}W_{R}^{3}+3V_{R}W_{L}^{2}W_{R}^{3}-4W_{L}^{3}W_{R}^{3}\big{)}$
$+$ $\left(-W_{L}V_{R}^{3}+W_{L}^{2}V_{R}^{2}-W_{L}^{3}V_{R}+W_{L}^{4}+V_{R}^{4}\right)$
×$\times\left(-V_{L}^{3}W_{R}+V_{L}^{2}W_{R}^{2}-V_{L}W_{R}^{3}+V_{L}^{4}+W_{R}^{4}\right).$
The factor $\left(V_{L}^{5}W_{L}^{5}-V_{R}^{5}W_{R}^{5}\right)$ is also present for the non-zero field case, and is solved by (B.3), however it does not lead to exceptional points. To analyze the factor $\left(\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}\right)U(u)$ we recall (III.25) for every possible quantized $k$ (five for $N=10$). For example, for $k=2\pi$,
$H_{1}(2\pi)H_{2}(2\pi)-u^{2}=\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}.$ (B.6)
Actually, we can check by brute force the following factorization,
$\left(\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}\right)U(u)=$
$\prod_{m=1}^{5}H_{1}(k_{m})H_{2}(k_{m})-u^{2}$ (B.7)
and confirm that the exceptional points give the roots of the discriminant.
B.2 Open boundary conditions
Now let us consider $N=11$, $\gamma=0$ (OBC) and $u=0$. In this case the discriminant (B.1) is given by,
$\text{disc}(P_{11}^{\text{(OBC)}}(z))=48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}f_{1}^{3}f_{2}^{3}f_{3}^{3}$ (B.8)
where
$f_{1}$ $=$ $V_{L}V_{R}+W_{L}W_{R}$
$f_{2}$ $=$ $V_{L}^{2}V_{R}^{2}-V_{L}V_{R}W_{L}W_{R}+W_{L}^{2}W_{R}^{2}$
$f_{3}$ $=$ $V_{L}^{2}V_{R}^{2}+V_{L}V_{R}W_{L}W_{R}+W_{L}^{2}W_{R}^{2}.$ (B.9)
Of course, the discriminant (B.8) vanishes when any of the couplings vanishes, consistent with our findings in the main text. The remaining factors can be verified by rewriting
$f_{1}f_{2}f_{3}=V_{L}^{5}V_{R}^{5}\sum_{j=0}^{5}\left(\frac{1}{\sqrt{\delta}}\right)^{2j}$ (B.10)
whose ten roots are given by (IV.16).
Next, we turn on the field $u$. In this case, we have,
$\text{disc}(P_{11}^{\text{(OBC)}}(z))$ $=$ $-48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}$ (B.11)
$\times$ $f_{1}^{2}f_{2}^{2}f_{3}^{2}g_{1}g_{2}g_{3}$
where $f_{j}$ is given by (B.2) and $g_{j}$ are given by
$g_{1}$ $=$ $u^{2}-(V_{L}V_{R}+W_{L}W_{R}),$
$g_{2}$ $=$ $u^{4}-2u^{2}f_{1}+f_{2},$
$g_{3}$ $=$ $u^{4}-2u^{2}f_{1}+f_{3}.$ (B.12)
Part of the roots remain the same as in the zero field case, but field dependent factors $g_{j}$ appear. These are associated with (IV.20). Indeed, by brute force computation we can verify
$\prod_{m=1}^{5}u^{2}-\tilde{H}_{1}(\theta_{m})\tilde{H}_{2}(\theta_{m})=g_{1}g_{2}g_{3}$ (B.13)
where $\theta_{m}$ is given by (IV.8). Let us remark that the product $g_{1}g_{2}g_{3}$ only depends on $\bar{u}$ and $\sqrt{\delta}$, up to an overall factor.
Finally, let us consider $N=12$ with non-zero field $u$. The discriminant is given by,
$\text{disc}(P_{12}^{\text{(OBC)}}(z))=4096V_{L}^{20}V_{R}^{20}W_{L}^{30}W_{R}^{30}U_{0}U_{1}(u)$ (B.14)
where
$U_{0}$ $=$ $153664V_{L}^{5}V_{R}^{5}+98000V_{L}^{4}V_{R}^{4}W_{L}W_{R}$ (B.15)
$+$ $60152V_{L}^{3}V_{R}^{3}W_{L}^{2}W_{R}^{2}+34709V_{L}^{2}V_{R}^{2}W_{L}^{3}W_{R}^{3}$
$+$ $17856V_{L}V_{R}W_{L}^{4}W_{R}^{4}+6912W_{L}^{5}W_{R}^{5}$
and
$U_{1}(u)$ $=$ $u^{12}-u^{10}(6V_{L}V_{R}+5W_{L}W_{R})$ (B.16)
$+$ $5u^{8}(3V_{L}^{2}V_{R}^{2}+4V_{L}V_{R}W_{L}W_{R}+2W_{L}^{2}W_{R}^{2})$
-- $2u^{6}(10V_{L}^{3}V_{R}^{3}+15V_{L}^{2}V_{R}^{2}W_{L}W_{R}+12V_{L}V_{R}W_{L}^{2}W_{R}^{2}$
$+5W_{L}^{3}W_{R}^{3})$
$+$ $u^{4}(15V_{L}^{4}V_{R}^{4}+20V_{L}^{3}V_{R}^{3}W_{L}W_{R}+18V_{L}^{2}V_{R}^{2}W_{L}^{2}W_{R}^{2}$
$+12V_{L}V_{R}W_{L}^{3}W_{R}^{3}+5W_{L}^{4}W_{R}^{4})$
-- $u^{2}(6V_{L}^{5}V_{R}^{5}+5V_{L}^{4}V_{R}^{4}W_{L}W_{R}+4V_{L}^{3}V_{R}^{3}W_{L}^{2}W_{R}^{2}$
$+3V_{L}^{2}V_{R}^{2}W_{L}^{3}W_{R}^{3}+2V_{L}V_{R}W_{L}^{4}W_{R}^{4}+W_{L}^{5}W_{R}^{5})$
$+$ $V_{L}^{6}V_{R}^{6}.$
Again, it is clear that the vanishing of any of the coupling parameters leads to a null discriminant (B.14). The factor $U_{0}$ can be written as
$U_{0}=153664V_{L}^{5}V_{R}^{5}\Big{(}1+\frac{125}{196}\frac{1}{\sqrt{\delta}^{2}}+\frac{7519}{19208}\frac{1}{\sqrt{\delta}^{4}}$
$+\frac{34709}{153664}\frac{1}{\sqrt{\delta}^{6}}+\frac{279}{2401}\frac{1}{\sqrt{\delta}^{8}}+\frac{108}{2041}\frac{1}{\sqrt{\delta}^{10}}\Big{)}.$ (B.17)
This factor is independent of the field $u$, and we can check numerically that the ten roots are given by (IV.25,IV.26), see summary in Table 1. The field dependent factor $U_{1}(u)$ can also be written in terms of $\sqrt{\delta}$ and $\bar{u}=u/(\sqrt{W_{L}}\sqrt{W_{R}})$, namely,
$U_{1}(u)=V_{L}^{6}V_{R}^{6}\Bigg{(}1-\frac{\bar{u}(3\bar{u}+1)}{\sqrt{\delta}^{2}}+\frac{\bar{u}(\bar{u}+1)(3\bar{u}^{2}-\bar{u}-1)}{\sqrt{\delta}^{4}}$
$-\frac{\bar{u}(\bar{u}-1)^{2}(\bar{u}+1)^{3}}{\sqrt{\delta}^{6}}\Bigg{)}\Bigg{(}1-\frac{\bar{u}(3\bar{u}-1)}{\sqrt{\delta}^{2}}$
$+\frac{\bar{u}(\bar{u}-1)(3\bar{u}^{2}+\bar{u}-1)}{\sqrt{\delta}^{4}}-\frac{\bar{u}(\bar{u}+1)^{2}(\bar{u}-1)^{3}}{\sqrt{\delta}^{6}}\Bigg{)}.$ (B.18)
We verify numerically that (IV.22,IV.23) do provide the roots of $U_{1}(u)=0$. In Table 2 we consider $\bar{u}=1/2$ and the associated solutions of (IV.22), as well as the associated exceptional $\sqrt{\delta}$’s.
Table 1: The ten roots of Eq. (B.2).
$\theta_{\text{EP}}^{\prime}$ $\sqrt{\delta}$
$05$ 0.209455 i $-0.628308+0.408561\text{i}$
$0.580505+0.209455\text{i}$ $-0.628308-0.408561\text{i}$
$1.07762-0.24362\text{i}$ $-0.344178+0.638086\text{i}$
$1.07762+0.24362\text{i}$ $-0.344178-0.638086\text{i}$
$1.5708-0.252951\text{i}$ $0.718376$
$1.5708+0.252951\text{i}$ $-0.718376$
$2.06398-0.24362\text{i}$ $0.344178+0.638086\text{i}$
$2.06398+0.24362\text{i}$ $0.344178-0.638086\text{i}$
$2.56109-0.209455\text{i}$ $0.628308+0.408561\text{i}$
$2.56109+0.209455\text{i}$ $0.628308-0.408561\text{i}$
Table 2: We consider $\bar{u}=0.5$ First six lines, $+$ solutions, last six lines -- solutions.
$\theta_{\text{EP}}^{\prime}$ $\sqrt{\delta}$
$0.642557-0.0955613\text{i}$ $-0.673088+0.294372\text{i}$
$0.642557+0.0955613\text{i}$ $-0.673088-0.294372\text{i}$
[…]
$1.5708+0.191123\text{i}$ $-0.694828\text{i}$
$2.49904-0.0955613\text{i}$ $0.673088+0.294372\text{i}$
$2.49904+0.0955613\text{i}$ $0.673088-0.294372\text{i}$
$0.343297$ $-1.31139$
$1.10745-0.172812\text{i}$ $-0.36085+0.604218\text{i}$
$1.10745+0.172812\text{i}$ $-0.36085-0.604218\text{i}$
$2.03414-0.172812\text{i}$ $0.36085+0.604218\text{i}$
$2.03414+0.172812\text{i}$ $0.36085-0.604218\text{i}$
$2.7983$ $1.31139$

### v3

@@
-Bulk-Boundary Correspondence and Exceptional Points for a Dimerized Hatano-Nelson Model with Staggered Potentials
+Exceptional Points, Bulk-Boundary Correspondence, and Entanglement Properties for a Dimerized Hatano-Nelson Model with Staggered Potentials
 Abstract
 It is well-known that the standard bulk-boundary correspondence does not hold for non-Hermitian systems in which also new phenomena such as exceptional points do occur.
-Here we study by analytical and numerical means a paradigmatic one-dimensional non-Hermitian model with dimerization, asymmetric hopping, and imaginary staggered potentials.
-We present analytical solutions for the eigenspectrum of this model with both open and closed boundary conditions as well as for the singular-value spectrum.
-We explicitly demonstrate the proper bulk-boundary correspondence between topological winding numbers in the periodic case and singular values in the open case.
+Here we study, mostly by analytical means, a paradigmatic one-dimensional non-Hermitian model with dimerization, asymmetric hopping, and imaginary staggered potentials.
+We present analytical solutions for the singular-value and the eigensystem of this model with both open and closed boundary conditions.
+We explicitly demonstrate that the proper bulk-boundary correspondence is between topological winding numbers in the periodic case and singular values, not eigenvalues, in the open case.
+These protected singular values are connected to hidden edge modes which only become exact zero-energy eigenmodes in the semi-infinite chain limit.
 We also show that a non-trivial topology leads to protected eigenvalues in the entanglement spectrum.
-In the $\mathcal{PT}$-symmetric case, we find that the model has a phase where exceptional points become dense in the thermodynamic limit.
+In the $\mathcal{PT}$-symmetric case, we find that the model has a so far overlooked phase where exceptional points become dense in the thermodynamic limit.
+This phase shows unusual hyper-ballistic transport properties with a dynamical critical exponent $z=1/2$.
 I Introduction
@@
 Strictly speaking, it requires a post-selection of those time evolutions which stay in the considered manifold [6].
-Non-Hermitian systems exhibit distinct properties associated with their complex spectrum and the difference between left and right eigenvectors [7].
+Another motivation to study non-Hermitian Hamiltonians came from realizing that hermiticity is not required to guarantee a real spectrum.
+The weaker condition of $\mathcal{PT}$-symmetry is sufficient, see e.g.
+Ref.
+[7] for a review.
+Generic non-Hermitian systems exhibit distinct properties associated with their complex spectrum and the difference between left and right eigenvectors [8].
 In this context, as with Hermitian systems, Gaussian non-Hermitian models can offer valuable insights into possible phases and phenomena.
-In particular, the topological properties of Gaussian non-Hermitian models have been studied and new phenomena have been revealed [8, 9, 10, 11, 12, 13, 11, 14, 15, 16].
+In particular, the topological properties of Gaussian non-Hermitian models have been studied and new phenomena have been revealed [9, 10, 11, 12, 13, 14, 15, 12, 16, 17, 18].
 For example, non-hermiticity leads to a proliferation of possible symmetry classes: the famous Altland-Zirnbauer ten-fold classification [19] for Hermitian Gaussian models is replaced by a 38-fold classification [9, 16] in the non-Hermitian domain.
 The richer classification scheme is a direct consequence of the distinction between transposition and conjugate transposition.
-Other distinctive properties include the extreme sensitivity of the eigenspectrum to the boundary conditions [15] leading to the skin effect and the breakdown of the conventional bulk-boundary correspondence [10, 12, 18, 19].
+Other distinctive properties include the extreme sensitivity of the eigenspectrum to the boundary conditions [13, 14, 17] leading to the skin effect and the breakdown of the conventional bulk-boundary correspondence [11, 14, 20, 21].
+As has very recently been proven, the proper bulk-boundary correspondence in non-Hermitian systems is based on the theory of Toeplitz operators and is between winding numbers of the symbols of such operators in the periodic case and the singular-value spectrum in the open case [22].
+Physically, it is important to note that the topologically protected singular values in general belong to boundary states which are metastable for finite systems and only become exact eigenstates in the limit of a semi-infinite system.
+The Hermitian case can then be understood as a special limit of this more general theory because in this case the singular values are just the absolute values of the eigenvalues.
 Another remarkable aspect of non-Hermitian Hamiltonians is the existence of so-called exceptional points [23] where both eigenvalues and eigenvectors coalesce.
 Such exceptional points can be detected in experiments [6].
-In order to highlight the different phenomena possible in Gaussian non-Hermitian models, we will study a minimal model which allows to turn various symmetries on and off.
+In order to highlight the different phenomena possible in Gaussian non-Hermitian models and to demonstrate in a fully analytical manner the recent novel results regarding their topology [22], we will study a minimal model which allows to turn various symmetries on and off.
 In the Hermitian case, one of the best-known one-dimensional systems with symmetry-protected topological order is the Su-Schrieffer-Heeger (SSH) model [24] which has chiral symmetry and belongs to the BDI class.
 This model has alternating strong and weak hoppings and topological edge modes are present if the two sites at the end of an open chain are weakly coupled to their respective neighbors.
-Here we make the model non-Hermitian by allowing for parity breaking hoppings, i.e., different hopping amplitudes to the left and to the right.
+Here, we make the model non-Hermitian by allowing for parity breaking hoppings, i.e., different hopping amplitudes to the left and to the right.
 In addition, we also introduce staggered imaginary onsite potentials to allow for a non-Hermitian, $\mathcal{PT}$-symmetric limit.
-Physically, the parity breaking can arise in an effective description of an SSH model with gain and loss.
+Physically, parity breaking can arise in an effective description of an SSH model with gain and loss.
 This phenomenon has been investigated in plasmonic chains [25] and photonic lattices [26].
-For equal hoppings to the left and to the right, the model is $\mathcal{PT}$ symmetric.
 The model is depicted in Fig.
@@
 […]
-Figure 1: The considered minimal model has a two-site unit cell, asymmetric intra- ($V_{L},V_{R}$) and inter-cell ($W_{L},W_{R}$) hopping as well as complex staggered potentials ±$\pm iu$.
+Figure 1: The considered minimal model has a two-site unit cell, asymmetric intra- ($V_{L},V_{R}$) and inter-cell ($W_{L},W_{R}$) hoppings as well as complex staggered potentials $\pm iu$.
 and can be viewed either as a dimerized Hatano-Nelson model [27] or an SSH chain with asymmetric hopping.
+The main motivation for our paper is twofold: (1) To demonstrate a mathematically rigorous bulk-boundary correspondence using the apparatus developed in Ref.
+[22] for a concrete, physically relevant model.
+(2) To re-examine the phase diagram of a model which contains many limiting cases which have been studied previously in the literature but to do so, as far as possible, analytically and in a rigorous manner instead of using exact diagonalizations.
+In the process, we discover a novel exceptional phase with unusual properties.
 Our paper is organized as follows: In Sec.
@@
 Furthermore, we determine the parameters for which the excitation gap closes and for which exceptional points do occur.
-We also introduce and determine the two winding numbers for this system.
+We also point out that exceptional points can occur in the complex plane of the model parameters.
+We introduce and determine the two winding numbers for this system.
+In the $\mathcal{PT}$-symmetric phase we discover a phase where exceptional points are dense leading to hyper-ballistic transport.
 In Sec.
@@
 This is also true for the conditions for the closing of the excitation gap and for the location of exceptional points.
+Depending on the parity of the length of the chain, we find different families of exceptional points which are determined via certain transcendental equations, also reflected in a factorization of the discriminant of the associated characteristic polynomial.
+They are clearly different from the closed boundary case.
 In Sec.
@@
 [22].
+We explicitly show in particular that the topologically protected boundary modes are metastable for finite system size and only become exact eigenstates in the limit of a semi-infinite system.
 In Sec.
-VI, we then discuss the entanglement entropy when cutting the system in half.
-The studied model is Gaussian and therefore its correlation matrix can be obtained [26, 27, 28, 29, 30].
+VI, we then discuss in detail the entanglement entropy when cutting the system in half.
+The studied model is Gaussian and therefore its correlation matrix can be obtained [28, 29, 30, 31, 32, 33, 34, 35].
 We find that in topological phases there are topologically protected eigenvalues in the entanglement spectrum.
+We note that the entanglement entropy is complex in general in the non-Hermitian case and find interesting properties of the entanglement entropy which can be directly tied to the physical properties of the system.
+In particular, we find that in the $\mathcal{PT}$-phase where exceptional points are dense the entanglement entropy is a nowhere continuous function.
 The last section is devoted to a short summary and our conclusions.
@@
 1.
-Complex staggered fields ±$\pm\text{i}u$ act on the sites of the chain.
+Complex staggered fields $\pm\text{i}u$ act on the sites of the chain.
 This model can be viewed either as a dimerized Hatano-Nelson model or as a Su-Schrieffer-Heeger (SSH) model with asymmetric hopping.
@@
 and $\bm{c}^{\dagger}=(c_{1}^{\dagger},\dots,c_{N}^{\dagger})$, $\bm{c}=(c_{1},\dots,c_{N})^{T}$ with $T$ denoting transposition.
+Note that in the semi-infinite case, the Hamilton operator has Toeplitz block form with $2\times 2$ matrices
+$h_{0}=\begin{pmatrix}iu&V_{L}\\ V_{R}&-iu\end{pmatrix}\;,\;h_{1}=\begin{pmatrix}0&0\\ W_{L}&0\end{pmatrix}\;,\;h_{-1}=\begin{pmatrix}0&W_{R}\\ 0&0\end{pmatrix}$ (II.5)
+which repeat along the diagonals. This will allow us later to define a corresponding symbol and winding numbers.
 To make this paper self-contained, we will recall the diagonalization of the asymmetric SSH chain for both closed ($\gamma=\pm 1)$ and open ($\gamma=0$) boundary conditions, referred to respectively as CBC or OBC.
 In both cases, the diagonalization of $\mathcal{T}_{N}$ with entries (II) follows from difference equations of the spectral problem.
-Here the lack of hermiticity of the single body Hamiltonian implies that left and right eigenvectors have to be considered.
+Here the lack of hermiticity of the single-body Hamiltonian implies that left and right eigenvectors have to be considered.
 The spectral problems read
-$\mathcal{T}_{N}\vec{r}=\epsilon~{}\vec{r},\quad(\mathcal{T}_{N})^{T}\vec{\ell}=\epsilon~{}\vec{\ell},$ (II.5)
-where $\vec{x}=(x_{1},\dots,x_{N})^{T}$ are right ($x=r$) and left ($x=\ell$) eigenvectors of $\mathcal{T}_{N}$, while $\epsilon$ is an eigenvalue.
+$\mathcal{T}_{N}~|\vec{r}\rangle=\epsilon~|\vec{r}\rangle,\quad\langle\vec{\ell}|~\mathcal{T}_{N}=\epsilon~\langle\vec{\ell}|,$ (II.6)
+where $|\vec{r}\rangle=(r_{1},\dots,r_{N})^{T}$ and $\langle\vec{\ell}|=(\ell_{1},\dots,\ell_{N})$ are right and left eigenvectors of $\mathcal{T}_{N}$, respectively, with eigenvalue $\epsilon$.
 The components of the right eigenvector satisfy the coupled difference equations
-$V_{L}r_{2j}+W_{R}r_{2j-2}=(\epsilon-\text{i}u)r_{2j-1},$ (II.6)
-$W_{L}r_{2j+1}+V_{R}r_{2j-1}=(\epsilon+\text{i}u)r_{2j},$ (II.7)
+$V_{L}r_{2j}+W_{R}r_{2j-2}=(\epsilon-\text{i}u)r_{2j-1},$ (II.7)
+$W_{L}r_{2j+1}+V_{R}r_{2j-1}=(\epsilon+\text{i}u)r_{2j},$ (II.8)
 with the boundary conditions
-$r_{0}=\gamma r_{N},\quad r_{N+1}=\gamma r_{1}.$ (II.8)
+$r_{0}=\gamma r_{N},\quad r_{N+1}=\gamma r_{1}.$ (II.9)
 Exchanging $R\leftrightarrow L$ one obtains the analogous equations for $\ell_{j}$, namely
-$V_{R}\ell_{2j}+W_{L}\ell_{2j-2}=(\epsilon-\text{i}u)\ell_{2j-1},$ (II.9)
-$W_{R}\ell_{2j+1}+V_{L}\ell_{2j-1}=(\epsilon+\text{i}u)\ell_{2j},$ (II.10)
+$V_{R}\ell_{2j}+W_{L}\ell_{2j-2}=(\epsilon-\text{i}u)\ell_{2j-1},$ (II.10)
+$W_{R}\ell_{2j+1}+V_{L}\ell_{2j-1}=(\epsilon+\text{i}u)\ell_{2j},$ (II.11)
 with the boundary conditions,
-$\ell_{0}=\gamma\ell_{N},\quad\ell_{N+1}=\gamma\ell_{1}.$ (II.11)
-For $\gamma=0$, we consider (II.8, II.11) for both parities of $N$, while for $\gamma\neq 0$ we only consider the case where $N$ is even.
+$\ell_{0}=\gamma\ell_{N},\quad\ell_{N+1}=\gamma\ell_{1}.$ (II.12)
+For $\gamma=0$, we consider (II.9, II.12) for both parities of $N$, while for $\gamma\neq 0$ we only consider the case where $N$ is even.
 In the following sections, each boundary condition is considered separately.
@@
 We only consider $N$ even for CBC.
-The difference equations (II.6,II.7,II.8) and (II.9,II.10,II.11) are easily solved by the ansatz
+The difference equations (II.7,II.8,II.9) and (II.10,II.11,II.12) are easily solved by the ansatz
 $r_{2j-1}=a_{R}(k)e^{-\text{i}kj},\quad r_{2j}=b_{R}(k)e^{-\text{i}kj},$ (III.1)
@@
 $H(k)|r^{\pm}(k)\rangle=\epsilon^{\pm}(k)|r^{\pm}(k)\rangle,$
-±$\langle\ell^{\pm}(k)|H(k)=\epsilon^{\pm}(k)\langle\ell^{\pm}(k)|,$ (III.6)
+$\langle\ell^{\pm}(k)|H(k)=\epsilon^{\pm}(k)\langle\ell^{\pm}(k)|,$ (III.6)
 where the quasienergies are given by
@@
 $=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{c}\text{i}u\pm\sqrt{H_{1}(k)H_{2}(k)-u^{2}}\\ H_{2}(k)\\ \end{array}\right),$ (III.13)
-±$\langle\ell^{\pm}(k)|$ $=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{cc}a_{L}^{\pm}(k),&b_{L}^{\pm}(k)\end{array}\right),$ (III.15)
+$\langle\ell^{\pm}(k)|$ $=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{cc}a_{L}^{\pm}(k),&b_{L}^{\pm}(k)\end{array}\right),$ (III.15)
 $=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{cc}\text{i}u\pm\sqrt{H_{1}(k)H_{2}(k)-u^{2}},&H_{1}(k)\,.\end{array}\right)$ (III.17)
@@
 $N^{\pm}=2\left(H_{1}(k)H_{2}(k)-u^{2}\pm\text{i}u\sqrt{H_{1}(k)H_{2}(k)-u^{2}}\right).$ (III.20)
-To conclude the solution, the parameter $k$ is quantized by (II.8) and given by
+To conclude the solution, the parameter $k$ is quantized by (II.9) and given by
 $k_{m}^{\text{(PBC)}}=\frac{2\pi(2m)}{N},\quad m=1,\dots,\frac{N}{2},$ (III.21)
@@
 $\Delta_{c}=(\epsilon^{+}(k)-\epsilon^{-}(k))/2=\sqrt{H_{1}(k)H_{2}(k)-u^{2}}.$ (III.23)
-The closing of both the real and the imaginary parts is well captured by the non-Hermitian gap [31]
-$\Delta=\text{min}_{k}|\Delta_{c}(k)|.$ (III.24)
-The complex gap (III.23) closes when,
+We define the non-Hermitian gap by [36]
+$\Delta=\text{min}_{k}|\Delta_{c}(k)|$ (III.24)
+which closes if
 $H_{1}(k)H_{2}(k)-u^{2}=0.$ (III.25)
@@
 Exceptional points and their relation with the discriminant of the characteristic polynomial of the hopping matrix $\mathcal{T}_{N}$ are discussed further in App.
-B.
-To analyze the manifold of exceptional points given by Eq.
+A.
+To analyze the phase transitions and the manifold of exceptional points given by Eq.
 (III.25) it is convenient to first set $u=0$.
@@
 Then, for every allowed quantized $k_{m}$ value and relaxing the requirement of real couplings, the solutions of $H_{1}=0$ or $H_{2}=0$ are
-$\left(\frac{V_{L}}{W_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}\quad\text{or}\quad\left(\frac{W_{L}}{V_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}.$ (III.26)
+$\left(\frac{V_{L}}{W_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}\quad\text{or}\quad\left(\frac{W_{L}}{V_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}$ (III.26)
 where $k_{m}$ is given by (III.21,III.22).
 It follows that for most values of $k_{m}$ the coupling parameters which lead to a gap closing/exceptional point are complex.
-For finite $N$ and periodic boundaries, one can see that $m=N/2$ ($k=2\pi$) and $m=N/4$ ($k=\pi$, if $N/2$ is even) are the only possible values of $k$ that leads to real ratios,
+For finite $N$ and periodic boundaries, one can see that $m=N/2$ ($k=2\pi$) and $m=N/4$ ($k=\pi$, if $N/2$ is even) are the only possible values of $k$ that lead to real ratios,
 $\left(\frac{V_{L}}{W_{R}}\right)_{\text{EP}}=\pm 1\quad\text{or}\quad\left(\frac{W_{L}}{V_{R}}\right)_{\text{EP}}=\pm 1.$ (III.27)
@@
 2.
-For finite $N$ and antiperiodic boundaries, the ratios (III.26) are real for $m=1/2(N/2-1)$ ($k=\pi$, if $N/2-1$ is even).
+They separate gapped phases labeled by two integer winding numbers $(\nu_{1},\nu_{2})$ whose definition is discussed in subsection III.3.
+With respect to the type of gap [16], the phases (0,1)(0,-1) and $(1,0)$ have a point gap while phases $(0,0)$ and (1,1)(1,-1) have a line gap (real or imaginary).
+For finite $N$ and antiperiodic boundaries, the ratios (III.26) are real for m=1/2(N/21)m=1/2(N/2-1) ($k=\pi$, if N/21N/2-1 is even).
 In the thermodynamic limit, the unit circles are filled and real ratios occur at $k=\pi$ or $k=2\pi$ for both periodic and antiperiodic boundary conditions.
-We remark that the sublattice-symmetric model has been widely considered in the literature, see e.g.
+We remark that sublattice-symmetric models have been widely considered in the literature, see e.g.
 Refs.
 [12, 15, 37, 38, 39, 40].
-Here we emphasize that the gap closing points and exceptional points coincide and that they can also occur in the complex parameter plane for every allowed, quantized value $k_{m}$.
-Figure 2: The non-Hermitian gap $\Delta$ as a function of the model parameters for $N=100$ and periodic boundary conditions.
-In panels (a) and (b) we set $V_{R}=W_{R}=1$ and $u=0$ and $u=0.5$ respectively.
-In panel (c) we consider $V_{L}=V_{R}=V$, $W_{L}=W_{R}=W$ and the field $u=1$.
-The numbers in panel (a) indicate the two winding numbers $(\nu_{1},\nu_{2})$, the numbers in panel (b) the winding number $I_{1}$.
-Now we consider a nonzero real field $u$.
+Here we emphasize that the gap closing points and exceptional points coincide and that they also occur in the complex parameter plane for every allowed quantized value $k_{m}$.
+Figure 2: The non-Hermitian gap $\Delta$, see Eq.
+(III.24), as a function of the model parameters in the thermodynamic limit for periodic boundary conditions.
+In panel (a), the zero field (sublattice symmetric) case, $u=0$, is shown for $V_{R}=W_{R}=1$ with two winding numbers $(\nu_{1},\nu_{2})$ characterizing each phase.
+In panel (b), we set $u=0.5$, therefore breaking the sublattice symmetry and leaving only a single winding number $I_{1}$ labeling the emerging phases.
+In panel (c), we consider the $\mathcal{PT}$-symmetric case $V_{L}=V_{R}=V$, $W_{L}=W_{R}=W$ and $u=1$.
+The regions in blue constitute a gapless phase where exceptional points are dense.
+The central “diamond” is a complex (anti-$\mathcal{PT}$) phase.
+The remaining phases are $\mathcal{PT}$-unbroken phases, either trivial or topological.
+Next, we consider a nonzero real field $u$.
 The field couples $H_{1}(k)$ and $H_{2}(k)$ and Eq.
@@
 2, the non-Hermitian gap (III.24) is shown for fixed parameters $V_{R}=W_{R}=1$ and $u=0.5$ as a function of $V_{L}$ and $W_{L}$.
-In this case, we can observe the merging of the zero field phases $(\nu_{1},\nu_{2})=(0,0)$ and $(1,-1)$ (we discuss the meaning of these numbers below) into a single phase as well as the bending of the phases boundaries.
-Another interesting case to consider is the $\mathcal{PT}$ symmetric case with $V_{L}=V_{R}=V$ and $W_{L}=W_{R}=W$ but the real field $u$ is arbitrary, such that $\sigma^{x}H(k)\sigma^{x}=H(k)^{*}$.
+The emerging phases are characterized by an integer winding number $I_{1}$ defined in subsection III.3.
+We can observe, for example, that the field $u$ induces the merging of the zero field phases $(\nu_{1},\nu_{2})=(0,0)$ and (1,1)(1,-1), with $V_{L}<-1$ and $W_{L}<-1$, into a single phase characterized by $I_{1}=0$.
+This can be understood by looking at the point $V_{L}=W_{L}=-1$, where $H_{1}(k)H_{2}(k)-u^{2}=-(2\sin(k/2)+\text{i}u)(2\sin(k/2)-\text{i}u)$ cannot be canceled for any real $k$ and $u$; therefore, the gap is open.
+On the other hand, the zero field phases $(0,0)$ and (1,1)(1,-1), with $V_{L}>1$ and $W_{L}>1$ do not merge at $u=0.5$, although the boundaries of the phases are bend.
+For instance, at $V_{L}=W_{L}=1$, one has $H_{1}(k)H_{2}(k)-u^{2}=(2\cos(k/2)-u)(2\cos(k/2)+u)$ which can be canceled by properly choosing a real $k$ depending on the value of the field, provided $|u|\leq 2$.
+Another interesting case to consider is the $\mathcal{PT}$ symmetric case with $V_{L}=V_{R}=V$ and $W_{L}=W_{R}=W$ and $u$ real but arbitrary, such that $\sigma^{x}H(k)\sigma^{x}=H(k)^{*}$.
 In this case, the constraint (III.25) reduces to,
@@
 2.
-The dark blue lines mark the the gap closing/ exceptional points, which only occur within the regions $|V-W|<1<|V+W|$ or $|V+W|<1<|V-W|$.
-In the thermodynamic limit, this region is gapless and filled with exceptional points.
-It thus constitutes an entire exceptional phase.
-We end this subsection remarking that in addition to the collision of quasienergies $\Delta_{c}=0$, another type of degeneracy in the spectrum can also be found.
-This degeneracy occurs within each band ϵ±superscriptitalic-$\epsilon^{\pm}$ for different quasimomentum $k$.
-Namely, one can impose
-$H_{1}(k)H_{2}(k)=H_{1}(k^{\prime})H_{2}(k^{\prime})$ (III.29)
-which gives another constraint on the model parameters.
-However, these degeneracies do not lead to exceptional points, as the eigenvectors do not coalesce, recall (III.13,III.17).
-This type of degeneracy, however, does appear as a factor in the discriminant of the characteristic polynomial of $\mathcal{T}_{N}$, see further remarks on this point in App.
-B, in particular Eq.
-(B.3).
+The blue regions |VW|<1<|V+W||V-W|<1<|V+W| or |V+W|<1<|VW||V+W|<1<|V-W| are gapless in the thermodynamic limit.
+This is a so far overlooked phase where the exceptional points are dense.
+In this phase, the $\mathcal{PT}$-symmetry is broken.
+The regions |V||W|>1|V|-|W|>1 and |W||V|>1|W|-|V|>1 are characterized by a purely real spectrum, and are adiabatically connected to the standard SSH model via $u\rightarrow 0$.
+Then, the former is called the $\mathcal{PT}$-unbroken trivial phase and the latter the $\mathcal{PT}$-unbroken topological phase.
+In the phase $|W|+|V|<1$ the spectrum is purely imaginary, and we name it the complex phase or anti-$\mathcal{PT}$ phase.
+The gapless phase where exceptional points are dense is peculiar because there is always a set of eigenvectors which is coalescent.
+I.e., this gapless phase is very different from a Hermitian phase where eigenvectors are orthogonal and one might wonder how this coalescence of eigenvectors affects its transport properties.
+To address this question, we investigate the scaling of the gap $\Delta\sim N^{-z}$ where $z$ is the dynamical critical exponent.
+A problem with such an analysis for a generic ratio of parameters $V/W$ is that the allowed momenta $4\pi m/N$ with $m=1,\cdots,N/2$ might or might not be commensurate or almost commensurate with the gap closing condition (III.28) depending on the system size $N$.
+In this case no clear scaling for generic $N$ emerges.
+To circumvent this issue, we choose a ratio $V/W$ which is badly approximable in the Diophantine sense by rational numbers thus making it possible to consider generic $N$.
+In Fig.
+3 we show, in particular, results for $V/W$ being the golden or the inverse golden ratio, respectively.
+Figure 3: Scaling of the gap (III.24) in the phase where exceptional points become dense for $V/W=(1-\sqrt{5})/2$ (black circles) and $V/W=(1+\sqrt{5})/2$ (red squares).
+The lines are fits $\Delta\sim N^{-1/2}$, i.e., the dynamical critical exponent is $z=1/2$.
+The results indicate that the dynamical critical exponent is $z=1/2$ which means that lengths scale as $\ell\sim t^{2}$.
+Transport in this phase is thus hyper-ballistic.
+We note that in a short-range Hermitian lattice model the Lieb-Robinson bounds make faster than ballistic transport impossible.
+However, such bounds do not apply in a non-Hermitian model where eigenvectors do not need to be orthogonal.
+In this phase, in particular, eigenvectors coalesce leading to an amplification.
 III.3 Winding numbers
-The eigenvalues of a non-Hermitian Hamiltonian are in general complex, and can trace a closed path around an arbitrary reference complex energy $E_{B}$, in which case a winding number can be assigned [36, 37, 18, 38, 32],
-[…]
-provided that $H(k)$ has a point gap, that is, $\det(H(k)-E_{B})\neq 0$.
+To characterize the different phases we have found, we can use topological invariants.
+The eigenvalues of a non-Hermitian Hamiltonian are in general complex and can trace, for periodic boundary conditions, a closed path around an arbitrary reference complex energy $E_{B}$, in which case a winding number can be assigned [41, 42, 20, 43, 37],
+$I_{1}=\frac{1}{2\pi\text{i}}\int_{0}^{2\pi}dk~\log\det(H(k)-E_{B}),$ (III.29)
+provided that $H(k)$ has a point gap at $E_{B}$, that is, $\det(H(k)-E_{B})\neq 0$.
 For concreteness we will chose $E_{B}=0$.
-Observe that the definition (III.30) is universal since no symmetry requirements are made.
+Observe that the definition (III.29) is universal since no symmetry requirements are made.
 Therefore, one dimensional non-Hermitian systems—in contrast to Hermitian ones—can have non-trivial topology in the absence of any spatial or non-spatial symmetries.
-If symmetries are present, then they lead to additional topological invariants [14].
+If symmetries are present, then they lead to additional topological invariants [16, 22].
 So far, mostly the case of a two-band model with sub-lattice symmetry has been considered [44, 45, 46, 47, 12, 43].
 The Hamiltonian of such a model takes the form
-$H(k)=\left(\begin{array}[]{cc}0&\hat{H}_{1}(k)\\ \hat{H}_{2}(k)&0\\ \end{array}\right)\,.$ (III.31)
+$H(k)=\left(\begin{array}[]{cc}0&\hat{H}_{1}(k)\\ \hat{H}_{2}(k)&0\\ \end{array}\right)\,.$ (III.30)
 Using the polar decomposition $\hat{H}_{j}(k)=d_{j}(k)q_{j}(k)$ with $q_{j}(k)$ unitary and $d_{j}(k)$ positive semi-definite, one can define two winding numbers
-[…]
+$\nu_{j}=\frac{1}{2\pi\text{i}}\int dk~{\rm tr}\,\left(q_{j}^{\dagger}\partial_{k}q_{j}\right)=\frac{1}{2\pi\text{i}}\int dk~\partial_{k}\log\det q_{j}\,.$ (III.31)
 Instead of the winding numbers $\nu_{1,2}$, one can equivalently consider [15, 48]
-$I_{1}=\nu_{1}+\nu_{2},$ (III.33)
-which is consistent with the definition of $I_{1}$ in Eq. (III.30), and
-$I_{2}=\nu_{2}-\nu_{1}$ (III.34)
-as independent winding numbers. For our model with $u=0$, we have $H_{j}(k)=|H_{j}(k)|e^{\text{i}\phi_{j}}$ and evaluating the integrals (III.33,III.34) leads to
-$I_{1}(u=0)=\begin{cases}-1,&\text{if}\quad~{}\left|\frac{W_{L}}{V_{R}}\right|>1\land\left|\frac{V_{L}}{W_{R}}\right|>1\\ 1,&\text{if}\quad\left|\frac{W_{L}}{V_{R}}\right|<1\land\left|\frac{V_{L}}{W_{R}}\right|<1\\ 0,&\text{otherwise}\end{cases}$ (III.35)
+$I_{1}=\nu_{1}+\nu_{2},$ (III.32)
+which is consistent with the definition of $I_{1}$ in Eq. (III.29), and
+$I_{2}=\nu_{2}-\nu_{1}$ (III.33)
+as independent winding numbers.
+Note that in the Hermitian case, $\hat{H}_{2}(k)=\hat{H}_{1}^{\dagger}(k)$, we have $\nu_{2}=-\nu_{1}$ implying that $I_{1}\equiv 0$ and only one non-trivial winding number, $I_{2}$, remains.
+For our model with $u=0$, we can write $H_{j}(k)=|H_{j}(k)|e^{\text{i}\phi_{j}}$ and evaluating the integrals (III.32,III.33) leads to
+$I_{1}(u=0)=\begin{cases}-1,&\text{if}\quad~\left|\frac{W_{L}}{V_{R}}\right|>1\land\left|\frac{V_{L}}{W_{R}}\right|>1\\ 1,&\text{if}\quad\left|\frac{W_{L}}{V_{R}}\right|<1\land\left|\frac{V_{L}}{W_{R}}\right|<1\\ 0,&\text{otherwise}\end{cases}$ (III.34)
 and
-$I_{2}=\begin{cases}-2&\text{if}\quad~{}\left|\frac{V_{L}}{W_{R}}\right|<1\land\left|\frac{W_{L}}{V_{R}}\right|>1\\ 0&\text{if}\quad\left|\frac{V_{L}}{W_{R}}\right|>1\land\left|\frac{W_{L}}{V_{R}}\right|<1\\ -1&\text{otherwise}.\end{cases}$ (III.36)
+$I_{2}=\begin{cases}-2&\text{if}\quad~\left|\frac{V_{L}}{W_{R}}\right|<1\land\left|\frac{W_{L}}{V_{R}}\right|>1\\ 0&\text{if}\quad\left|\frac{V_{L}}{W_{R}}\right|>1\land\left|\frac{W_{L}}{V_{R}}\right|<1\\ -1&\text{otherwise}.\end{cases}$ (III.35)
 Clearly, the winding numbers are not defined when $\left|\frac{V_{L}}{W_{R}}\right|=\left|\frac{W_{L}}{V_{R}}\right|=1$ which are precisely the gapless/exceptional lines discussed in the previous subsection.
-The two winding numbers $\nu_{1,2}$ are shown in the left panel of Fig.
-2 for parameters $V_{R}=W_{R}=1$ (no loss of generality) in each phase as a function of $V_{L}$ and $W_{L}$.
-Next, we briefly discuss the $\mathcal{PT}$-symmetric case.
-The Hamiltonian (III.3) for $V_{L}=V_{R}=V$ and $W_{L}=W_{R}=W$ satisfies the $\mathcal{PT}$ symmetry $\sigma^{x}H(k)\sigma^{x}=H(k)^{*}$ as well as chiral symmetry $\sigma^{z}H(k)^{\dagger}\sigma^{z}=-H(k)$, and this case is part of the AZ symmetry class AIII [14].
-The phase diagram for $u=1$ is shown in Fig.
-2(c) and consists of gapped phases and phases where exceptional points become dense in the thermodynamic limit.
-These phases will be discussed further in Sec.
-VI when we analyze the entanglement spectrum.
+The two winding numbers $\nu_{1,2}=(I_{1}\mp I_{2})/2$ are shown in the left panel of Fig.
+2 in each phase as a function of $V_{L}$ and $W_{L}$.
 IV Open boundary conditions
@@
 In App.
-A, we provide some details on both approaches.
+B, we provide some details on both approaches.
 The connection between these two approaches was pointed out in Ref.
@@
 [52, 53].
-The solution of the open boundary case can also be obtained by a simple transformation to the standard (symmetric) SSH chain [12], namely,
-$S\mathcal{T}_{N}S^{-1}=\mathcal{T}_{N}^{SSH}$ (IV.1)
+The solution of the open boundary case can also be obtained by a simple similarity transformation to the standard (symmetric) SSH chain which leaves the spectrum invariant [14], namely,
+$S^{-1}\mathcal{T}_{N}S=\mathcal{T}_{N}^{SSH}$ (IV.1)
 where $S$ is a diagonal matrix with entries
-$s_{2j-1}=\frac{V_{R}^{\frac{j-1}{2}}W_{R}^{\frac{j-1}{2}}}{V_{L}^{\frac{j-1}{2}}W_{L}^{\frac{j-1}{2}}},\quad s_{2j}=\frac{V_{R}^{\frac{j}{2}}W_{R}^{\frac{j}{2}}}{V_{L}^{\frac{j}{2}}W_{L}^{\frac{j}{2}}}\frac{W_{L}}{W_{R}},$ (IV.2)
+$s_{2j-1}=\frac{V_{R}^{\frac{j-1}{2}}W_{R}^{\frac{j-1}{2}}}{V_{L}^{\frac{j-1}{2}}W_{L}^{\frac{j-1}{2}}},\quad s_{2j}=\frac{V_{R}^{\frac{j}{2}}W_{R}^{\frac{j}{2}}}{V_{L}^{\frac{j}{2}}W_{L}^{\frac{j}{2}}}\frac{W_{L}^{\frac{1}{2}}}{W_{R}^{\frac{1}{2}}},$ (IV.2)
 and $\mathcal{T}_{N}^{SSH}$ is a symmetric hopping matrix of the standard SSH chain with effective couplings,
-$V_{\text{eff}}=\sqrt{V_{L}V_{R}},\quad W_{\text{eff}}=\sqrt{W_{L}W_{R}}.$ (IV.3)
+$V_{\text{eff}}=\sqrt{V_{L}}\sqrt{V_{R}},\quad W_{\text{eff}}=\sqrt{W_{L}}\sqrt{W_{R}}.$ (IV.3)
 The quasienergies are then given by
@@
 $\theta_{m}^{(OBC)}=\frac{2\pi m}{N+1},\quad m=1,\dots,\frac{N-1}{2},$ (IV.8)
-which gives $N-1$ eigenvalues from (IV.4) and (IV.8). Together with the isolated root $\epsilon=\text{i}u$, one has all the $N$ eigenvalues of the tridiagonal matrix when $\gamma=0$.
+which gives N1N-1 eigenvalues from (IV.4) and (IV.8). Together with the isolated root $\epsilon=\text{i}u$, one has all the $N$ eigenvalues of the tridiagonal matrix when $\gamma=0$.
 For even $N$, the parameter $\theta$ is determined via the transcendental equation
 $\sqrt{\delta}\sin\left(\left(\frac{N}{2}+1\right)\theta\right)+\sin\left(\frac{N}{2}\theta\right)=0\,.$ (IV.9)
-We remark that the transcendental equation (IV.9) has the same form as the transcendental equation of the quantum Ising chain with transverse field $\lambda=\sqrt{\delta}$ and $L/2$ sites [49].
+We remark that the transcendental equation (IV.9) has the same form as the transcendental equation of the quantum Ising chain with transverse field $\lambda=\sqrt{\delta}$ and $N/2$ sites [54].
 Meanwhile, the eigenvector components can be written as
@@
 $\left(\sqrt{\delta}\right)_{\text{EP}}=-e^{-\text{i}\theta_{m}}~~\text{or}~~\left(\sqrt{\delta}\right)_{\text{EP}}=-e^{\text{i}\theta_{m}}.$ (IV.16)
-Therefore, for non-vanishing hopping parameters, there are $N-1$ exceptional points.
+Therefore, for non-vanishing hopping parameters, there are N1N-1 exceptional points associated with the eigenvalue $\epsilon=0$.
 For finite $N$, these points always have an imaginary part.
@@
 $\tilde{H}_{1}(\theta_{m})\tilde{H}_{2}(\theta_{m})-u^{2}=\sqrt{V_{L}}\sqrt{V_{R}}\sqrt{W_{L}}\sqrt{W_{R}}$
-×$\quad\times\left(\sqrt{\delta}+\frac{1}{\sqrt{\delta}}(1-\bar{u}^{2})+2\cos(\theta_{m})\right)\,.$ (IV.19)
-The minimal gap (IV.15) as a function of $\bar{u}$ and $\delta$ is shown in Fig. 3.
-Figure 3: The rescaled minimal gap $\Delta$ as a function of $\sqrt{\delta}$ and $\bar{u}$ for $N=101$ and open boundary conditions.
-The exceptional points are therefore now given by
+$\quad\times\left(\sqrt{\delta}+\frac{1}{\sqrt{\delta}}(1-\bar{u}^{2})+2\cos(\theta_{m})\right)\,.$ (IV.19)
+The exceptional points, which are still equivalent to the points where the gap closes, are therefore now given by
 $\left(\sqrt{\delta}\right)_{\text{EP}}=-\cos(\theta_{m})\pm\sqrt{\bar{u}^{2}-\sin^{2}(\theta_{m})}.$ (IV.20)
-Interestingly, the expression (IV.20) implies that real exceptional points do exist for finite $N$ provided that $\bar{u}\geq|\sin(\theta_{m})|$ or $\bar{u}\leq-|\sin(\theta_{m})|$.
-In the thermodynamic limit, critical fields $\bar{u}\rightarrow\pm 1$ emerge when $\theta_{m}$ approaches 00 or $\pi$.
+Interestingly, the expression (IV.20) implies that now real exceptional points do exist even for finite $N$ provided that $\bar{u}\geq|\sin(\theta_{m})|$ or $\bar{u}\leq-|\sin(\theta_{m})|$.
+In the thermodynamic limit, critical fields $\bar{u}\rightarrow\pm 1$ emerge when $\theta_{m}$ approaches 0 or $\pi$.
 In Fig.
 4 some examples of the loci of exceptional points for different values of the rescaled field $\bar{u}$ are shown.
-Figure 4: Loci of exceptional points $\left(\sqrt{\delta}\right)_{\text{EP}}$ for $N=41$ and different rescaled fields $\bar{u}$.
-Recall that when $N$ is odd, there is, in addition, an isolated root $\epsilon=\text{i}u$.
+Figure 4: Loci of exceptional points $\left(\sqrt{\delta}\right)_{\text{EP}}$, see Eq. (IV.20), for $N=41$ and different rescaled fields $\bar{u}$.
+Recall that when NN is odd, there is, in addition, an isolated root $\epsilon=\text{i}u$.
 We can impose that this root coincides with the quasienergies (IV.4), that is, $\text{i}u=\epsilon^{\pm}$.
@@
 $\sqrt{\delta}+\frac{1}{\sqrt{\delta}}(1-\bar{u}^{2})+2\cos(\theta)=0$ (IV.21)
-but with $\theta$ now given by by (IV.9).
+but with $\theta$ now given by (IV.9).
 Since Eq.
@@
 The transcendental equation (IV.25) is clearly independent of the field $u$.
-It can be solved numerically; there are $N-2$ solutions [50] with $\text{Re}(\theta_{\text{EP}}^{\prime})\geq 0$ which fully describe the field independent part of the discriminant.
+It can be solved numerically; there are N2N-2 solutions [55] with $\text{Re}(\theta_{\text{EP}}^{\prime})\geq 0$ which fully describe the field independent part of the discriminant.
 The numerical solution for different lattice sizes is shown in the upper panel of Fig.
@@
 V Bulk-boundary correspondence
-It is well-known that the standard bulk-boundary correspondence for topological Hermitian systems breaks down in the non-Hermitian case.
-Mathematically, the fundamental issue is that the eigenvalues of a finite non-Hermitian matrix do, in general, not converge to the eigenvalues of the semi-infinite case.
+For Hermitian systems, the bulk-boundary correspondence connects a topological invariant calculated for a system with periodic boundary conditions with the number of topologically protected edge modes in a system with open boundaries which have energies which converge exponentially to zero with system size.
+It is well-known that this type of bulk-boundary correspondence breaks down in the non-Hermitian case [16, 14, 21].
+Some authors have tried to devise new topological invariants based on, for example, ’non-Bloch theory’ [14] but these attempts lack a solid mathematical foundation.
+Furthermore, these theories are based on translational invariance, however, ”topological features must be properties of a translationally not invariant Hamiltonian” [56] and thus should be stable against perturbations which break translational invariance.
+From a mathematical point of view, none of this is at all surprising: it is well-known from Toeplitz theory that topological invariants are in general related to properties of the singular value, not the eigenvalue, spectrum.
+The importance of the singular value spectrum in the non-Hermitian case has been pointed out, for example, already in Refs.
+[37, 36] but without establishing a rigorous bulk-boundary correspondence.
+An analysis of the singular value spectrum has also turned out to be useful in the classification of non-Hermitian random matrices [57].
+It is also worth noting that another way of understanding the importance of singular values is Hermitization [58] where the hermitian block matrix
+$\bar{H}=\begin{pmatrix}0&H\\ H^{\dagger}&0\end{pmatrix}$
+is constructed.
+A point gap in $H$ then becomes a spectral gap in $\bar{H}$ and the positive eigenvalues of $\bar{H}$ correspond to the singular values of $H$.
+However, this approach does introduce an unphysical additional chiral symmetry and is not needed here.
+The case where the Hamiltonian $H$ itself is Hermitian is special because the singular values are then just the absolute values of the eigenvalues so that index theorems for singular values also apply to the eigenvalues.
+Another important and so far overlooked issue is that the eigenvalues of a finite non-Hermitian matrix do, in general, not converge to the eigenvalues of the semi-infinite case.
 This is similar to a discontinuous point.
-More specifically, in the Hermitian case we can connect the topological invariant calculated for a periodic system with the number of edge states in a finite system.
-The energy of these edge states converges exponentially to zero with system size.
-This is what is known as the standard bulk-boundary correspondence.
-In the non-Hermitian case, the eigenspectrum cannot be used to construct a bulk boundary correspondence for a finite system.
-Instead, it is known from Toeplitz theory that the singular value spectrum has to be used.
-The K-splitting theorem states [51] that a system with winding number $I$ has $K\geq|I|$ singular values which will converge to zero with increasing system size and which belong to states which are exponentially localized at the boundary.
-In the semi-infinite case these states become exact eigenstates with exactly zero energy but for finite system size they only get mapped exponentially close to zero by the Hamiltonian but are not exact eigenstates.
-I.e., in a finite non-Hermitian system the zero energy edge modes are ’hidden’, in general, and constitute very long-lived states.
-The proper bulk-boundary correspondence for non-Hermitian systems was recently fully explored in Ref.
+All of these aspects can be fully understood if one starts from known index theorems in Toeplitz theory.
+In particular, the K-splitting theorem states [59] that a system with winding number $I$ has $K\geq|I|$ singular values which will converge to zero with increasing system size and which belong to states which are exponentially localized at the boundary.
+These topologically protected singular values are separated by a gap from the rest of the singular value spectrum.
+In the semi-infinite case, these states become exact eigenstates with exactly zero energy but for finite system size they only get mapped exponentially close to zero by the Hamiltonian but are not exact eigenstates.
+I.e., in a finite non-Hermitian system the topological zero-energy edge modes are ’hidden’, in general, and instead constitute very long-lived metastable states.
+The latter aspect is extremely important for a proper understanding of the topological properties of non-Hermitian systems and for testing these properties experimentally.
+One possible scenario for such an experiment is to prepare a system in a hidden zero mode—which is localized at one of the boundaries—in a cold atomic gas, trapped ion system, or some other quantum simulation/computation platform and to monitor a generalized Loschmidt echo.
+This Loschmidt echo will stay close to $1$ for a time which is proportional to system size thus indicating the metastable character and its convergence to a true eigenstate in the thermodynamic limit.
+The corresponding protocol is discussed in more detail in the Suppl.
+Mat.
+of Ref.
 [22].
-We note that this formalism does encompass the Hermitian case because for a Hermitian matrix the singular values are just the absolute values of the eigenvalues.
-Here we want to demonstrate the bulk-boundary correspondence explicitly for the model under consideration.
+We note that the required postselection to trajectories without jumps has already been performed for a single qubit [6].
+Such an experiment, generalized to multiple qubits, would be able to detect the hidden zero modes.
+We stress again that the formalism developed in Ref.
+[22] does encompass the standard bulk-boundary correspondence in the Hermitian case because for a Hermitian matrix the singular values are just the absolute values of the eigenvalues.
+Here we want to explicitly and fully analytically demonstrate this general bulk-boundary correspondence for the first time for the physically relevant minimal non-Hermitian model under consideration.
 V.1 Singular values
@@
 $s_{2j-2}+s_{2j+2}=\frac{\sigma^{2}-(W_{L}^{2}+V_{R}^{2})}{W_{L}V_{R}}s_{2j}$ (V.3)
-with boundaries fixed by the last line in (V.1). Using the method reviewed in App. A, we find that there are two solutions to the difference equations (V.1).
+with boundaries fixed by the last line in (V.1). Using the method reviewed in App. B, we find that there are two solutions to the difference equations (V.1).
 In the first (second) solution the even (odd) components of the vector $\vec{s}$ are zero $s_{2j}=0$ ($s_{2j-1}=0$) and the left boundary fixes the ratio $s_{3}/s_{1}$ ($s_{4}/s_{2}$).
@@
 The spectrum of (V.1) is given by the union of $\sigma_{\phi}^{2}$ and $\sigma_{\omega}^{2}$.
-Remarkably, the pairs of variables $\{V_{L},W_{R}\}$ and $\{V_{R},W_{L}\}$ decouple from each other.
+Remarkably, the pairs of variables $\{V_{L},W_{R}\}$ and $\{V_{R},W_{L}\}$ decouple from each other showing the value of an analytical solution.
 This means that the characteristic polynomial associated with (V.1) factorizes into two independent factors depending solely on each of these pairs.
@@
 This complex root produces an exponentially small singular value $\sigma_{\phi}$.
-Similarly, we find that (V.11) has $N/2$ real solutions in the interval $0<\omega<\pi$ when $|V_{R}/W_{L}|\geq 1$ and $N/2-1$ real solutions plus a complex conjugate pair solution when $|V_{R}/W_{L}|<1$.
+Similarly, we find that (V.11) has $N/2$ real solutions in the interval $0<\omega<\pi$ when $|V_{R}/W_{L}|\geq 1$ and N/21N/2-1 real solutions plus a complex conjugate pair solution when $|V_{R}/W_{L}|<1$.
 In Fig.
@@
 The Bloch Hamiltonian for the corresponding periodic problem with $u=0$ has block structure, see Eq.
+(III.30).
+There are therefore two winding numbers $\nu_{1,2}$ defined in Eq.
 (III.31).
-There are therefore two winding numbers $\nu_{1,2}$ defined in Eq.
-(III.32).
 The K-splitting theorem then predicts that there are $K\geq|\nu_{1}|+|\nu_{2}|$ singular values $\sigma$ with $\lim_{N\to\infty}\sigma=0$ [59, 22].
-This is fully consistent with our results, compare Fig.
+This is fully consistent with our results; compare Fig.
 7 with the phase diagram in the left panel of Fig.
@@
 For a finite system, the singular values have to be considered as we have done above.
-Here we want to show explicitly how the exact edge states in the semi-infinite limit arise and that they become ’hidden’ if the system is finite.
+Here we want to show explicitly how the exact edge states in the semi-infinite limit arise and how they become ’hidden’ if the system is finite.
 For a semi-infinite chain, we can find zero modes by looking for solutions of $Hv=0$ and $\tilde{H}w=0$.
@@
 For the odd vector components we find the recurrence relation $V_{R}v_{2j-1}+W_{L}v_{2j+1}=0$ with the solution $v_{2j+1}=\left(-\frac{V_{R}}{W_{L}}\right)^{j}v_{1}$.
-In addition, we have to demand that the solution is normalizable, $||v||^{2}=|v_{1}|^{2}\sum_{j=0}^{\infty}\left(-\frac{V_{R}}{W_{L}}\right)^{2j}=1$.
+In addition, we have to demand that the solution is normalizable, $v||^{2}=|v_{1}|^{2}\sum_{j=0}^{\infty}\left(-\frac{V_{R}}{W_{L}}\right)^{2j}=1$.
 This is possible if $|W_{L}|>|V_{R}|$.
@@
 They are then no longer eigenvectors, $Hv\neq 0$, but they do get mapped exponentially close to zero, i.e.
-$||Hv||\to 0$ for $N\to\infty$ and similarly for $w$.
+$Hv||\to 0$ for $N\to\infty$ and similarly for $w$.
 This is what we mean by hidden zero modes.
 For the example considered in Fig.
-8 with $V_{R}=W_{R}=1$ this means that the first zero mode is stable for $-1<V_{L}<1$ and the second zero mode for $|W_{L}|>1$.
+8 with $V_{R}=W_{R}=1$ this means that the first zero mode is stable for $1<V_{L}<1$ and the second zero mode for $|W_{L}|>1$.
 This is fully consistent with the singular values as well as with the winding numbers as is expected based on the K-splitting theorem.
@@
 The results show that the singular values are stable and thus indeed topologically protected while the eigenvalues are, in general, not protected.
-More specifically, we find that for the case $W_{L}=0.5$ shown in the lower left panel there are two degenerate zero eigenvalues in the unperturbed system for $-0.4\lesssim V_{L}\lesssim 0.5$.
+More specifically, we find that for the case $W_{L}=0.5$ shown in the lower left panel there are two degenerate zero eigenvalues in the unperturbed system for $0.4\lesssim V_{L}\lesssim 0.5$.
 However, the small perturbation immediately moves them away from zero energy.
@@
 8.
-In the unperturbed case, there are two degenerate zero energy eigenvalues for $-1.3\lesssim V_{L}\lesssim 1.5$.
+In the unperturbed case, there are two degenerate zero energy eigenvalues for $1.3\lesssim V_{L}\lesssim 1.5$.
 Interestingly, adding a perturbation does not completely remove them from zero energy.
-Instead, they are stable for $-1<V_{L}<1$.
+Instead, they are stable for $1<V_{L}<1$.
 This can be understood as follows: in this regime, the non-Hermitian Hamiltonian with $V_{R}=W_{R}=1$ and $W_{L}=1.5$ can be adiabatically connected to a chiral Hermitian Hamiltonian without closing the gap.
@@
 VI Entanglement entropy
-In this section we investigate the entanglement entropy for the non-Hermitian model (II) with periodic boundary conditions.
-We use the reduced density matrix approach for free fermions [52, 53] which has been extended to the non-Hermitian realm [27, 28, 54, 26, 55] using the generalized density matrix $\rho=|r\rangle\langle\ell|$ where $\langle\ell|$ and $|r\rangle$ are bi-orthogonal left and right eigenstates of (II).
+For a system in a topological phase, we expect that a non-zero lower bound for the entanglement entropy exists.
+This entanglement is protected and cannot be removed [60].
+To study this aspect of topology, we investigate in this section the entanglement entropy for the non-Hermitian model (II) with periodic boundary conditions.
+The other novel aspect which motivates our study of the entanglement entropy is the $\mathcal{PT}$-symmetric phase in which exceptional points are dense.
+As we will show, this leads to an entanglement entropy which is nowhere continuous.
+We use the reduced density matrix approach for free fermions [61, 62] which has been extended to the non-Hermitian realm [28, 29, 30, 31, 32, 33, 34, 35, 63] using the generalized density matrix $\rho=|r\rangle\langle\ell|$ where $\langle\ell|$ and $|r\rangle$ are bi-orthogonal left and right eigenstates of (II).
 The key quantity is the two-point correlation matrix with elements
-$C_{ij}=\langle\ell|c_{i}^{\dagger}c_{j}|r\rangle={\rm tr}\,\left(\rho c_{i}^{\dagger}c_{j}\right)$ (VI.1)
+$C_{ij}=\langle\ell|c_{i}^{\dagger}c_{j}|r\rangle={\rm tr}\,\left(\rho c_{i}^{\dagger}c_{j}\right)\,.$ (VI.1)
 Restricting the indices to a given subsystem $\mathcal{A}$, the von Neumann entanglement entropy is given by,
@@
 In this paper we only consider $\mathcal{A}=\{1,\dots,N/2\}$, that is, we cut the system in half.
-For the model considered in this paper, in contrast to Hermitian systems, the eigenvalues $\nu_{j}$ and $S_{\mathcal{A}}$ can be complex.
+For the model considered in this paper, in contrast to Hermitian systems, the eigenvalues $\nu_{j}$ and therefore also $S_{\mathcal{A}}$ can be complex.
 Nevertheless, we will indiscriminately call $S_{\mathcal{A}}$ the entanglement entropy.
 In addition, we choose the branch cut of the logarithm to be on the negative real axis.
+Before we proceed further, we might ask why $S_{\mathcal{A}}$ is complex and what the meaning of a complex entanglement entropy is.
+For the first part of the question it is important to remember that a non-Hermitian Hamiltonian is an effective approach obtained from a Master equation by dropping terms which induce quantum jumps.
+The density matrix in the proper Master equation is always positive semi-definite.
+$S_{\mathcal{A}}$ being complex is an artefact of the approximation.
+Nevertheless, the entanglement entropy can be a useful indicator for the correlations in the system and we expect, in particular, that it is sensitive to phase transitions and exceptional points.
+Physically, one can typically understand the real part of $S_{\mathcal{A}}$ as being still associated with the spreading of Schmidt values while the imaginary part is related to dissipation.
 Also, in non-Hermitian systems, the notion of a ground state is not well defined since the spectrum is complex in general.
@@
 We now investigate the spectrum of the correlation matrix (VI.12) and the associated entanglement entropy for indices in (VI.12) restricted to $i,j\in\{1,\dots,N/2\}$.
-We perform numerical calculations for large lattices, in which case we only keep eigenvalues of the correlation matrix that satisfy $|\nu_{j}|>10^{-50}$ and $|1-\nu_{j}|>10^{-50}$.
-Additionally, we set to zero real or imaginary parts of $\nu_{j}$ and $1-\nu_{j}$ that are individually smaller than $10^{-50}$.
-This is particular important in cases where the eigenvalues of the correlation matrix are negative with a tiny imaginary part, whose sign can numerically oscillate around the branch cut.
-Before considering numerics, as a warm-up, we study the trivial 2-cell model ($N=4$) analytically.
+We perform numerical calculations for large lattices, in which case we only keep eigenvalues of the correlation matrix that satisfy $|\nu_{j}|>\Lambda$ and $|1-\nu_{j}|>\Lambda$.
+Additionally, we set to zero real or imaginary parts of $\nu_{j}$ and $1-\nu_{j}$ that are individually smaller in magnitude than the chosen cutoff $\Lambda=10^{-50}$.
+This is particularly important in cases where the eigenvalues of the correlation matrix are negative with a tiny imaginary part, whose sign can numerically oscillate around the branch cut.
+By calculating the discarded weight $\sum_{|\nu_{j}|<\Lambda}|\nu_{j}|+\sum_{|1-\nu_{j}|<\Lambda}|1-\nu_{j}|$ we make sure that the cutoff does not affect the results on the scale shown in the figures provided that we stay on one branch of the logarithm.
+Before considering numerical results for large systems we will also study, as a warm-up, the 2-cell model ($N=4$) analytically.
 Both zero field $u=0$ and $\mathcal{PT}$-symmetric cases are considered.
 VI.1 Sub-lattice symmetric case
-For the zero field case, setting $V_{R}=W_{R}=1$, let us consider transitions between the phases $(\nu_{1},\nu_{2})=(1,-1)$ (with $W_{L}>1$) and $(0,-1)$, between $(1,0)$ and $(0,0)$ and between $(1,-1)$ (with $W_{L}<-1$) and $(0,-1)$, recall the phase diagram shown in the left panel of Fig.
+For the zero field case, setting $V_{R}=W_{R}=1$, let us consider transitions between the phases (i) $(\nu_{1},\nu_{2})=(1,-1)$ (with $W_{L}>1$) and (0,1)(0,-1), (ii) between $(1,0)$ and $(0,0)$, and (iii) between (1,1)(1,-1) (with $W_{L}<-1$) and (0,1)(0,-1), recall the phase diagram shown in the left panel of Fig.
 2.
 For concreteness, let us consider only $V_{L}\geq 0$.
-Recall that the spectrum of $\mathcal{T}_{N}$ in the phases $(1,-1)$ and $(0,0)$ is in general composed of two lobes in the complex plane either separated by the imaginary axis or real axis, therefore being characterized by $I_{1}=0$.
-On the other hand, the spectrum in the phases $(1,0)$ and $(0,-1)$ traces a closed path around $E_{B}=0$, therefore being characterized by $I_{1}\neq 0$.
+Recall that the spectrum of $\mathcal{T}_{N}$ in the phases (1,1)(1,-1) and $(0,0)$ is in general composed of two lobes in the complex plane either separated by the imaginary axis or real axis, therefore being characterized by $I_{1}=0$.
+On the other hand, the spectrum in the phases $(1,0)$ and (0,1)(0,-1) traces a closed path around $E_{B}=0$, therefore being characterized by $I_{1}\neq 0$.
 We start with the 2-cell model ($N=4$).
 For this lattice size, there is no lobe in the spectrum, and the eigenvalues are either purely real or purely imaginary.
-We fill the states with energy $-\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$ and leave empty the states with energy $\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$.
+We fill the states with energy $\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$ and leave empty the states with energy $\sqrt{(V_{L}\mp 1)(1\mp W_{L})}$.
 The associated correlation matrix is given by,
@@
 9 for some fixed values of $W_{L}$ as function of $V_{L}$.
-We observe that within the phases characterized by $(1,-1)$ the eigenvalues form conjugate pairs of the form $1/2\pm\text{i}\alpha$, see panels (a) and (c) in Fig.
+We observe that within the phases characterized by (1,1)(1,-1) the eigenvalues form conjugate pairs of the form $1/2\pm\text{i}\alpha$, where $\alpha$ is a given function of $V_{L}$ and $W_{L}$, see panels (a) and (c) in Fig.
 9 for the fixed values $W_{L}=\pm 1.5$.
@@
 On the other hand, deep inside the phase, taking $V_{L}=0$ and $W_{L}\rightarrow\infty$, we have $\alpha\rightarrow 0$ and therefore $S_{2}\rightarrow 2\log 2$.
-Increasing $V_{L}$ and crossing the transition point $V_{L}=1$, we observe that the eigenvalues of the correlation matrix in the phase $(0,-1)$ are complex, with the form $\beta_{1,2}\pm\text{i}\alpha$.
+Increasing $V_{L}$ and crossing the transition point $V_{L}=1$, we observe that the eigenvalues of the correlation matrix in the phase (0,1)(0,-1) are complex, with the form $\beta_{1,2}\pm\text{i}\alpha$.
 That is, the eigenvalues have imaginary parts with opposite signs but with different real parts $\beta_{2}=1-\beta_{1}$.
@@
 However, for large $|W_{L}|$ the imaginary part tends to disappear when $V_{L}$ increases.
-In the limit $V_{L}\rightarrow\infty$ and $|W_{L}|\rightarrow\infty$, the exact eigenvalues of the correlation matrix are $41/2\pm\sqrt{2}/$ 4, producing a real entropy $S_{2}\approx 0.832991$.
+In the limit $V_{L}\rightarrow\infty$ and $|W_{L}|\rightarrow\infty$, the exact eigenvalues of the correlation matrix are $1/2\pm\sqrt{2}/4$, producing a real entropy $S_{2}\approx 0.832991$.
 In the phase $(1,0)$, the eigenvalues once more have the form $\beta_{1,2}\pm\text{i}\alpha$ with $\beta_{2}=1-\beta_{1}$, and the entanglement entropy is complex (except if $V_{L}=W_{L}$), see panel (b) in Fig.
 9 for the fixed value $W_{L}=0.1$.
-Moving to the phase $(0,0)$, we observe that the eigenvalues of $C$ are real, however, they are greater than unity or smaller then zero, and therefore also produce a complex entanglement entropy.
-Only in the limit $V_{L}\rightarrow\infty$ and $W_{L}\rightarrow 0$, do the eigenvalues of the correlation matrix tend to $1$ and 00, thus producing a zero entanglement entropy.
-Figure 9: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the trivial two-cell model ($N=4$) with zero field $u=0$, $W_{R}=V_{R}=1$ and some fixed values of $W_{L}$ indicated in the labels (a)-(c) as a function of $V_{L}$.
+Moving to the phase $(0,0)$, we observe that the eigenvalues of $C$ are real, however, they are greater than unity or smaller than zero, and therefore also produce a complex entanglement entropy.
+Only in the limit $V_{L}\rightarrow\infty$ and $W_{L}\rightarrow 0$, do the eigenvalues of the correlation matrix tend to $1$ and 0, thus producing a zero entanglement entropy.
+Figure 9: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the two-cell model ($N=4$) with zero field $u=0$, $W_{R}=V_{R}=1$ and some fixed values of $W_{L}$ indicated in the labels (a)-(c) as a function of $V_{L}$.
+$S_{2}$ is discontinuous at the phase transitions.
 Now, we consider larger lattice sizes using numerical diagonalization.
-We observe a variety of possible types of eigenvalues in the correlation spectrum, similar to the trivial 2-cell model.
+We observe a variety of possible types of eigenvalues in the correlation spectrum, similar to the 2-cell model.
 In some phases, we may find eigenvalues with $\text{Re}(\nu_{j})>1$ as well as $\text{Im}(\nu_{j})\neq 0$.
-Also, we note that near the transition points the behaviour of the correlation matrix eigenvalues is more intricate.
-Let us analyze each phase in details.
+Also, we note that near the transition points the behavior of the correlation matrix eigenvalues is more intricate.
+Let us now analyze each phase in detail.
 Results are collected in Fig.
@@
 The lattice size is $N=120$.
+$S_{N/2}$ is discontinuous at the phase transitions.
 In the insets, the typical form of the complex spectrum is shown; the occupied states are indicated by filled circles while the unoccupied states by open circles.
@@
 For better visualization, in the insets we consider $N=52$.
-[…]
-Moving to the phases $(0,-1)$, see once more panels (a) and (c) in Fig.
-10, we note that all eigenvalues of the correlation matrix satisfy $10<\text{Re}(\nu_{j})<$ 1, with most eigenvalues being either close to unity or zero.
+Starting with the phases (1,1)(-1,1), see panels (a) and (c) in Fig.
+10 for the fixed values $W_{L}=\pm 1.5$, we observe that most of the eigenvalues of the correlation matrix are either close to $1$ or 0.
+Some of the eigenvalues have the property $\nu_{j}>1$ or $\nu_{j}<0$, and there are two eigenvalues sitting at $1/2\pm\text{i}\alpha$.
+As the lattice size increases, we observe that those eigenvalues with $\nu_{j}>1$ or $\nu_{j}<0$ tend to unity or zero, while $\alpha$ decreases.
+For instance, for $N=52$, the lattice size considered in the insets, the highest eigenvalue is $\nu_{\text{max}}\approx 1.000000000290214$, the lowest $\nu_{\text{min}}=1-\nu_{\text{max}}$ and $\alpha\approx 4.275663\times 10^{-4}$, while for $N=120$ the highest eigenvalue for $W_{L}=1.5$ and $V_{L}=0.5$ is $\nu_{\text{max}}\approx 1.00000000000000000265940$, the lowest $\nu_{\text{min}}=1-\nu_{\text{max}}$ and $\alpha\approx 2.955440\times 10^{-7}$.
+Therefore, the entanglement entropy comes essentially from the conjugate eigenvalues $1/2\pm\text{i}\alpha$, and its value is also given by (VI.14).
+For a fixed length of the chain, but going deeper into the phase by increasing $|W_{L}|$, we also note that $\alpha$ decreases.
+Then, in the limit $\alpha\rightarrow 0$, the entanglement entropy has the finite value $S_{N/2}\rightarrow 2\log 2$ deep inside the phase.
+Similar considerations can be made for the phase (1,1)(1,-1) with $W_{L}<-1$; note however that a different filling must be used in this case.
+Moving to the phases (0,1)(0,-1), see once more panels (a) and (c) in Fig.
+10, we note that all eigenvalues of the correlation matrix satisfy $0<\text{Re}(\nu_{j})<1$, with most eigenvalues being either close to unity or zero.
 Some intermediate eigenvalues appear, and these carry a small imaginary part.
@@
 It follows that, remarkably, this phase may be described by a CFT with central charge $c=1$, even though it has a point gapped spectrum.
-Similar results can be obtained for $W_{L}<-1$ in the same phase $(0,-1)$.
-Figure 11: Scaling of the entanglement entropy in the phase $(0,-1)$ for $W_{L}=1.5$ and values of $V_{L}$ deep inside the phase.
+Similar results can be obtained for $W_{L}<-1$ in the same phase (0,1)(0,-1).
+Figure 11: Scaling of the entanglement entropy in the phase (0,1)(0,-1) for $W_{L}=1.5$ and values of $V_{L}$ deep inside the phase.
 The phase $(1,0)$, see panel (b) in Fig.
-10, is similar to the phases $(0,-1)$.
+10, is similar to the phases (0,1)(0,-1).
 The scaling of the entanglement entropy for this case is shown in Fig.
@@
 Then, four regions must be analyzed.
-The region $|W-V|<1<W+V$ is the exceptional phase, characterized by a large number of exceptional points (recall Sec.
-III.2).
-In this phase, the $\mathcal{PT}$-symmetry is broken.
-The regions $W-V<-1$ and $W-V>1$ are adiabatically connected to the standard SSH model via $u\rightarrow 0$.
-Then, the former is called the $\mathcal{PT}$-unbroken trivial phase and the latter the $\mathcal{PT}$-unbroken topological phase.
-In the phase $W+V<1$ the spectrum is purely imaginary, and we name it the complex phase.
-As before, we start our analysis by the simple 2-cell model ($N=4$), which has the following correlation matrix,
+As before, we start our analysis with the simple 2-cell model ($N=4$), which has the following correlation matrix,
 $C=\frac{1}{4}\left(\begin{array}[]{cc}\frac{(V-W)^{2}}{(V-W)^{2}-1+\text{i}\sqrt{(V-W)^{2}-1}}+\frac{(V+W)^{2}}{(V+W)^{2}-1+\text{i}\sqrt{(V+W)^{2}-1}}&\frac{W-V}{\sqrt{(V-W)^{2}-1}}-\frac{V+W}{\sqrt{(V+W)^{2}-1}}\\ \frac{W-V}{\sqrt{(V-W)^{2}-1}}-\frac{V+W}{\sqrt{(V+W)^{2}-1}}&\frac{(V-W)^{2}}{(V-W)^{2}-i\sqrt{(V-W)^{2}-1}-1}+\frac{(V+W)^{2}}{(V+W)^{2}-i\sqrt{(V+W)^{2}-1}-1}\\ \end{array}\right)$
-associated with the filled states $-\sqrt{(V\mp W)^{2}-1}$ and empty states $\sqrt{(V\mp W)^{2}-1}$.
+associated with the filled states $\sqrt{(V\mp W)^{2}-1}$ and empty states $\sqrt{(V\mp W)^{2}-1}$.
 We plot in Fig.
 13 the eigenvalues of $C$ and the associated entanglement entropy for some fixed values of $V$ as a function of $W$, covering all the four regions of the phase diagram.
-For $V=0$ and $10<W<$ 1 that is, within the complex phase, see panel (a) in Fig.
-13, we can observe that the eigenvalues of the correlation matrix are real with $\nu_{j}>1$ or $\nu_{j}<0$ until the phase transition point $W=1$ is reached.
+For $V=0$ and $0<W<1$ that is, within the complex phase, see panel (a) in Fig.
+13, we observe that the eigenvalues of the correlation matrix are real with $\nu_{j}>1$ or $\nu_{j}<0$ until the phase transition point $W=1$ is reached.
 As a consequence, the entanglement entropy is complex in this phase.
@@
 Increasing $V$ to $V=0.5$, see panel (b) in Fig.
-13, the behaviour of the eigenvalues $\nu_{j}$ is the same as for $V=0$ until the exceptional phase is reached at $W=0.5$.
+13, the behavior of the eigenvalues $\nu_{j}$ is the same as for $V=0$ until the exceptional phase is reached at $W=0.5$.
 Within the exceptional phase, the eigenvalues $\nu_{j}$ have the form $\beta_{1,2}\pm\text{i}\alpha$, that is, the eigenvalues have the same imaginary part with opposite signs but different real parts related by $\beta_{2}=1-\beta_{1}$.
@@
 To conclude the analysis of the 2-cell model, we consider $V=1.5$ such that the $\mathcal{PT}$-unbroken trivial phase is considered.
-This phase for $V=1.5$ occurs when $0.50<W<$ 0.5, and we note that there the eigenvalues are real with $\nu_{j}>1$ or $\nu_{j}<0$, producing again a complex entanglement entropy except if $W=0$.
-However, deep inside the phase ($V\rightarrow\infty$), the eigenvalues tend to 00 and $1$, leading to a null entanglement, as expected from a trivial phase.
-For $2.50.5<W<$ 2.5 we encounter once more the exceptional phase and for $W>2.5$ the $\mathcal{PT}$-unbroken topological phase.
-Figure 13: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the trivial $\mathcal{PT}$-symmetric two-cell model ($N=4$) with $u=1$, some fixed values of $V$ indicated in the labels (a)-(c) as a function of $W$.
+This phase for $V=1.5$ occurs when $0<W<0.5$, and we note that there the eigenvalues are real with $\nu_{j}>1$ or $\nu_{j}<0$, producing again a complex entanglement entropy except if $W=0$.
+However, deep inside the phase ($V\rightarrow\infty$), the eigenvalues tend to 0 and $1$, leading to a vanishing entanglement entropy, as expected from a trivial phase.
+For $0.5<W<2.5$ we encounter once more the exceptional phase and for $W>2.5$ the $\mathcal{PT}$-unbroken topological phase.
+Figure 13: Eigenvalues of the correlation matrix (top and middle panels) and entanglement entropy (bottom panels) for the $\mathcal{PT}$-symmetric two-cell model ($N=4$) with $u=1$, and fixed values of $V$ as indicated in the labels (a)-(c) as a function of $W$.
+$S_{2}$ is discontinuous at the phase transitions.
 Finally, we consider the $\mathcal{PT}$ case for large lattice sizes.
@@
 The same four regions considered for the 2-cell model are analyzed.
-First, let us consider $V=0$ and vary $W$.
-For $10<W<$ 1, we note that the correlation matrix has real eigenvalues, most of them 00 or $1$, but there is one eigenvalue $\nu_{\text{max}}>1$ and one eigenvalue $\nu_{\text{min}}<0$.
-Remarkably, these eigenvalues do not depend on the lattice size, and therefore correspond to the two eigenvalues of the 2-cell model.
-For instance, for $V=0,W=0.5$, the eigenvalues are $\nu_{\text{max}}\approx 1.077350$ and $\nu_{\text{min}}\approx-0.077350$, which produces a finite entanglement entropy with negative real part and positive imaginary part, precisely the same as the 2-cell model, recall panel (a) in Fig.
-13.
-Crossing the transition point $W=1$, we note that most of the eigenvalues of the correlation matrix are 00 or $1$, but there are a couple of eigenvalues of the form $1/2\pm\text{i}\alpha$, where $\alpha$ also does not depend on the lattice size.
-However, we note that for a fixed length $\alpha\rightarrow 0$ when $W\rightarrow\infty$.
-Therefore, like in some of the previously considered cases, the entanglement approaches $S_{N/2}\rightarrow 2\log(2)$ deep inside this phase.
-Now we consider $V\neq 0$.
-Similarly to the 2-cell model, by fixing $10<V<$ 1, we can reach the complex phase, the exceptional phase and the $\mathcal{PT}$-unbroken topological phase by varying $W$.
+Similarly to the 2-cell model, by fixing $0<V<1$, we can reach the complex phase, the exceptional phase and the $\mathcal{PT}$-unbroken topological phase by varying $W$.
 On the other hand, by fixing $V>1$ and varying $W$ we can reach the $\mathcal{PT}$-unbroken trivial phase, the exceptional phase and the $\mathcal{PT}$-unbroken topological phase.
 Let us consider once more $V=0.5$.
-Within the complex phase ($0.50<W<$ 0.5 for $V=0.5$), all the eigenvalues of the correlation matrix jump above unity or below zero, see panel (a) in Fig.
+Within the complex phase ($0<W<0.5$ for $V=0.5$), all the eigenvalues of the correlation matrix jump above unity or below zero, see panel (a) in Fig.
 14 where the parameters $W=0.45$ and $V=0.5$ are used.
@@
 15.
-We note that two eigenvalues stand out, being significantly greater than $1$ or less than 00.
+We note that two eigenvalues stand out, being significantly greater than $1$ or less than 0.
 All the eigenvalues of the correlation matrix in this phase seem to vary slowly with the size of the chain, actually tending to stabilize.
@@
 Also, the phase is surrounded by critical lines that drive the entanglement entropy to high values.
-Moving to the $\mathcal{PT}$-unbroken topological phase, we note that the correlation spectrum is composed by eigenvalues close to $1$ and 00, with some of them being slightly bigger than $1$ or smaller then 00, as well as a conjugate pair of the form $1/2\pm\text{i}\alpha$, see panel (b) in Fig.
+Moving to the $\mathcal{PT}$-unbroken topological phase, we note that the correlation spectrum is composed by eigenvalues close to $1$ and 0, with some of them being slightly bigger than $1$ or smaller then 0, as well as a conjugate pair of the form $1/2\pm\text{i}\alpha$, see panel (b) in Fig.
 14 where the parameters are set to $W=1.7,V=0.5$.
-As the lattice size increases, we note that the eigenvalues tend to $1$ or 00, while $\alpha$ decreases slowly and seems to stabilize with the length of the chain.
+As the lattice size increases, we note that the eigenvalues tend to $1$ or 0, while $\alpha$ decreases slowly and seems to stabilize with the length of the chain.
 The associated entanglement entropy is real and finite.
@@
 15.
-In the $\mathcal{PT}$-unbroken trivial phase, the correlation spectrum is also composed mostly of 00’s and $1$’s, some of them slightly bigger than $1$ or smaller than 00.
+In the $\mathcal{PT}$-unbroken trivial phase, the correlation spectrum is also composed mostly of 0’s and $1$’s, some of them slightly bigger than $1$ or smaller than 0.
 Also, some eigenvalues carry an imaginary part in the form of conjugate pairs, see panel (d) in Fig.
@@
 Interestingly, similarly to previous cases, both real and imaginary part of the eigenvalues vary slowly with the length of the chain.
-Nevertheless, deep inside this phase (large $V$), the eigenvalues tend to 00 or $1$ and produces zero entanglement, as expected from a trivial phase; see panel (b) in Fig.
+Nevertheless, deep inside this phase (large $V$), the eigenvalues tend to 0 or $1$ and produces zero entanglement, as expected from a trivial phase; see panel (b) in Fig.
 15.
-Let us briefly discuss the exceptional phase.
-In this phase we observe an intricate behaviour of the correlation spectrum, characterized by various complex eigenvalues, see panel (c) in Fig.
-14 where the parameters are set to $V=0.5,\,W=1.2$.
-This behaviour is reflected in an intricate evolution of the entanglement entropy, see panels (a) and (b) in Fig.
+Let us now discuss the exceptional phase.
+In this phase we observe an intricate behavior of the correlation spectrum, characterized by various complex eigenvalues, see panel (c) in Fig.
+14 where the parameters are set to V=0.5,W=1.2V=0.5,\,W=1.2.
+This behavior is reflected in an intricate evolution of the entanglement entropy, see panels (a) and (b) in Fig.
 15.
@@
 This is numerically challenging and goes beyond the scope of this paper.
-In brief, the entanglement entropy has a rich behaviour in the $\mathcal{PT}$ symmetric model and is capable of detecting exceptional points and phase transitions.
+In brief, the entanglement entropy has a rich behavior in the $\mathcal{PT}$ symmetric model and is capable of detecting both exceptional points and phase transitions.
 In the $\mathcal{PT}$-unbroken topological phase there are two protected eigenvalues at $1/2$, contributing $2\log 2$ to the entanglement entropy.
@@
 In the insets, we zoom into regions where the exceptional points are very close to each other.
-VII Conclusion
+In the thermodynamic limit, the exceptional points become close and $S_{N/2}$ is nowhere continuous.
+VII Conclusions
 In this paper we have analytically studied a paradigmatic one-dimensional, non-Hermitian lattice model.
-Our goal was to illustrate many of the unique features of non-Hermitian systems in a concrete model and to do so largely based on analytical instead of purely numerical results.
+Our goal was to illustrate many of the unique features of non-Hermitian systems—in particular, the proper bulk-boundary correspondence and the exceptional points—in a concrete model and to do so largely based on analytical instead of purely numerical results.
+A particular focus of our work was to clarify the recently described relation between topological winding numbers for periodic systems and topologically protected singular values as well as hidden zero modes in open systems [22].
+We have shown explicitly that the latter only become exact zero-energy eigenmodes in the limit of a semi-infinite chain.
+For a finite chain, they correspond to metastable states instead which is a crucial distinction for experiments.
 We derived the eigensystems both for closed (periodic and anti-periodic) as well as open boundary conditions.
@@
 In the considered model, we have shown that such points typically occur when the complex gap closes and that the location of exceptional points, in general, differs between the closed and the open case.
-Surprisingly, we found in the $\mathcal{PT}$-symmetric phase that exceptional points can even become dense in the thermodynamic limit, forming an entire “exceptional phase”.
-One of the central results of this work is an explicit demonstration of a proper bulk-boundary correspondence for a non-Hermitian system.
+Surprisingly, we found in the $\mathcal{PT}$-symmetric case that exceptional points can even become dense in the thermodynamic limit, forming an entire novel “exceptional phase”.
+This phase has also other unique properties which are impossible in a Hermitian phase.
+We found, in particular, that transport is hyper-ballistic with a dynamical critical exponent $z=1/2$.
+As already mentioned above, one of the central results of this work is an explicit demonstration of a proper bulk-boundary correspondence for a non-Hermitian system.
 This point has lead to some confusion in the literature.
@@
 Furthermore, we demonstrated that in topologically non-trivial phases there are protected eigenvalues in the entanglement spectrum which lead to lower, non-trivial entanglement bounds.
-Non-trivial topology in the studied non-Hermitian system thus leads to two observable and experimentally relevant phenomena: (1) Extremely long-lived stable zero energy states, and (2) non-trivial entanglement in systems with a gap between occupied and unoccupied energy bands.
-Appendix A Some details for the open boundary case
-To make this paper self-contained, we review some details of the diagonalization of $\mathcal{T}_{N}$ (II) for open boundary conditions, that is, $\gamma=0$, following [44, 45].
-The approach in [44] is based on recurrence relations satisfied by the characteristic polynomial,
-$P_{N}(z)=\det_{N}\left(\mathcal{T}_{N}-z\right).$ (A.1)
-The roots $P_{N}(z)=0$ are the quasienergies of the Hamiltonian (II.1). The characteristic polynomial (A.1) for $\gamma=0$ is special as it satisfies the recurrence relations
-$P_{2j}(z)$ $=$ $(-\text{i}u-z)P_{2j-1}(z)-V_{L}V_{R}P_{2j-2}(z),$
-$P_{2j-1}(z)$ $=$ $(\text{i}u-z)P_{2j-2}(z)-W_{L}W_{R}P_{2j-3}(z)$ (A.2)
-depending on the parity of $N$. Note that one has an additional recurrence relation for the characteristic polynomial, namely
-$P_{j}(z)$ $=$ $\left((z+\text{i}u)(z-\text{i}u)-(V_{L}V_{R}+W_{L}W_{R})\right)P_{j-2}(z)$ (A.3)
--- $V_{L}V_{R}W_{L}W_{R}P_{j-4}(z),$
-where appropriate initial conditions must be chosen.
-On the other hand, following [45], one notes that equations (II.6,II.7) can be iterated to the same fourth order difference equation,
-$r_{j+2}+\frac{V_{R}W_{R}}{V_{L}W_{L}}r_{j-2}$
-$=\left(-\frac{(\text{i}u+\epsilon)(\text{i}u-\epsilon)+(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}}\right)r_{j}.$ (A.4)
-The difference equation (A) can be compared to (A.3) leading to the solution (IV.10).
-In addition, following [45], introduce
-$T_{n}(\alpha,\beta)=\frac{\alpha^{n+1}-\beta^{n+1}}{\alpha-\beta}$ (A.5)
-and the sequence
-$a_{n+2}-(\alpha+\beta)a_{n+1}+\alpha\beta a_{n}=0.$ (A.6)
-Then, it follows that
-$a_{n}=T_{n-2}(\alpha,\beta)a_{2}-\alpha\beta T_{n-3}(\alpha,\beta)a_{1}$ (A.7)
-is the solution of the difference equation (A.6). Now choosing
-$\alpha+\beta=\frac{(\text{i}u-\epsilon)(-\text{i}u-\epsilon)-(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}},$ (A.8)
-$\alpha\beta=\frac{V_{R}W_{R}}{V_{L}W_{L}},$ (A.9)
-it follows that
-$r_{2j}=T_{j-2}(\alpha,\beta)r_{4}-\alpha\beta T_{j-3}(\alpha,\beta)r_{2},$ (A.10)
-$r_{2j-1}=T_{j-2}(\alpha,\beta)r_{3}-\alpha\beta T_{j-3}(\alpha,\beta)r_{1},$ (A.11)
-is a solution of (II.6,II.7,A). The boundary conditions determine,
-$r_{1}$ $=$ $V_{L},\quad r_{2}=\epsilon-\text{i}u,\quad r_{3}=W_{R}+V_{L}(\alpha+\beta),$
-$r_{4}$ $=$ $(\epsilon-\text{i}u)\left(2\frac{W_{R}}{V_{L}}+\alpha+\beta\right).$ (A.12)
-where $r_{1}=V_{L}$ is an arbitrary choice. The conditions (A.8) give $\{\epsilon,\beta\}$ in terms of $\alpha$,
-$\epsilon$ $=$ ±$\pm\sqrt{V_{L}V_{R}+W_{L}W_{R}+V_{L}W_{L}(\alpha+\beta)-u^{2}},$
-$\beta$ $=$ $\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}.$ (A.13)
-Using the parametrization
-$\alpha=\frac{\sqrt{V_{R}}\sqrt{W_{R}}}{\sqrt{V_{L}}\sqrt{W_{L}}}\exp(\text{i}\theta),$ (A.14)
-and then
-$T_{n}\left(\alpha,\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}\right)=\frac{V_{R}^{\frac{n}{2}}W_{R}^{\frac{n}{2}}}{V_{L}^{\frac{n}{2}}W_{L}^{\frac{n}{2}}}\frac{\sin((n+1)\theta)}{\sin(\theta)}$ (A.15)
-in (A.10,A.11) we find (IV.10,LABEL:vLodd) with the initial conditions (A).
-The boundary condition at the end of the chain gives a condition on the parameter $\theta$, which depends on whether $N$ is odd or even.
-For odd $N$, the condition $x_{L+1}=0$ comes from (IV.10) with $12n=L+$ 1 and gives the quantized solution (IV.8).
-For even $N$, the condition $x_{L+1}=0$ comes from (LABEL:vLodd) with $12n-1=L+$ 1 and leads to the transcendental equation (IV.9).
-The solution (IV.12,LABEL:vLoddL) is obtained analogously.
-Appendix B Discriminant of the characteristic polynomial
+A non-trivial topology in non-Hermitian systems thus leads to two observable and experimentally relevant phenomena: (1) Extremely long-lived metastable zero energy states, and (2) non-trivial entanglement in systems with a gap between occupied and unoccupied energy bands.
+We end this section by mentioning some possible directions for future studies.
+First, it would be worthwhile to study the loci of exceptional points for other families of non-Hermitian free-fermionic models, including free fermions and parafermions in disguise [64, 65].
+More importantly, it would be very interesting to extend these studies to the loci of exceptional points for non-Hermitian interacting models, see e.g.
+Ref.
+[66] for inspiration.
+Second, it would be important to go back to other models which have been studied in the literature using incorrect ideas of topology and to work out the proper bulk-boundary relations.
+For one specific example this has already been done in Ref.
+[22].
+Finally, the scaling behavior of the entanglement entropy at the various critical lines observed in this work certainly deserves further investigations.
+The exceptional phase in the $\cal{PT}$-symmetric case is particularly intriguing, and studying this novel phase further might advance our understanding of critical behavior in non-Hermitian systems.
+Acknowledgments
+R.A.P.
+acknowledges support from the Rio Grande do Sul State Research Support Foundation (FAPERGS) and the National Council for Scientific and Technological Development (CNPq).
+J.S.
+acknowledges support by NSERC via the Discovery grants program.
+Appendix A Discriminant of the characteristic polynomial
 We recall that a polynomial has repeated roots if and only if its discriminant
-$\text{disc}(P_{N}(z))=\text{disc}\left(\det_{N}\left(\mathcal{T}_{N}-z\right)\right)$ (B.1)
+$\text{disc}(P_{N}(z))=\text{disc}\left(\det_{N}\left(\mathcal{T}_{N}-z\right)\right)$ (A.1)
 vanishes.
 The discriminant is given by the determinant of the Sylvester matrix [67], and it is in general an intricate expression of the parameters of the model.
-The vanishing of the discriminant (B.1) is a necessary condition for the existence of an exceptional point.
+The vanishing of the discriminant (A.1) is a necessary condition for the existence of an exceptional point.
 Of course, the discriminant vanishes when the quasienergies collide.
@@
 In the main text, the conditions for gap closing/exceptional points are determined from the simple characterization of the eigenvalues by (III.7) and (IV.4).
-Here these results are verified by explicitly computing the discriminant (B.1) using Mathematica for different boundary conditions.
-B.1 Closed boundary conditions
-Let us consider $\gamma=1$ and $u=0$. For a lattice with $N=10$ sites the discriminant (B.1) is given by,
-$\text{disc}(P_{10}^{\text{(PBC)}}(z))$ $=$ $10^{10}\left(W_{L}^{5}+V_{R}^{5}\right)\left(V_{L}^{5}+W_{R}^{5}\right)$ (B.2)
+Here these results are verified by explicitly computing the discriminant (A.1) using Mathematica for different boundary conditions.
+A.1 Closed boundary conditions
+Let us consider $\gamma=1$ and $u=0$. For a lattice with $N=10$ sites the discriminant (A.1) is given by,
+$\text{disc}(P_{10}^{\text{(PBC)}}(z))$ $=$ $10^{10}\left(W_{L}^{5}+V_{R}^{5}\right)\left(V_{L}^{5}+W_{R}^{5}\right)$ (A.2)
 $\times$ $\left(V_{L}^{5}W_{L}^{5}-V_{R}^{5}W_{R}^{5}\right)^{8}.$
 The roots of the factors $\left(W_{L}^{5}+V_{R}^{5}\right)\left(V_{L}^{5}+W_{R}^{5}\right)$ are clearly given by (III.26).
-The extra factor comes from (III.29), which can be seen by writing (III.29) like,
-$\frac{V_{L}W_{L}}{V_{R}W_{R}}=e^{\text{i}(k+k^{\prime})}=e^{\frac{2}{5}\text{i}\pi(\ell+\ell^{\prime})}$ (B.3)
+The extra factor comes from another type of degeneracy in the spectrum which can occur within each band $\epsilon^{\pm}$ for different quasimomentum $k$.
+Namely, one can impose
+$H_{1}(k)H_{2}(k)=H_{1}(k^{\prime})H_{2}(k^{\prime})$ (A.3)
+which gives another constraint on the model parameters.
+Note that these degeneracies do not lead to exceptional points, as the eigenvectors do not coalesce, recall (III.13,III.17).
+The extra factor in the discriminant is indeed given by (A.3) which vanishes for
+$\frac{V_{L}W_{L}}{V_{R}W_{R}}=e^{\text{i}(k+k^{\prime})}=e^{\frac{2}{5}\text{i}\pi(\ell+\ell^{\prime})}$ (A.4)
 with $\ell+\ell^{\prime}=1,\dots,5$.
-As mentioned in the main text, this last factor produce degenerate eigenvalues of $\mathcal{T}_{N}$ but associated with different eigenvectors, thus they are not exceptional points.
 Now let us consider again $N=10$, $\gamma=1$, but with $u\neq 0$. In this case the discriminant is given by,
@@
 $+$ $\left(-W_{L}V_{R}^{3}+W_{L}^{2}V_{R}^{2}-W_{L}^{3}V_{R}+W_{L}^{4}+V_{R}^{4}\right)$
-×$\times\left(-V_{L}^{3}W_{R}+V_{L}^{2}W_{R}^{2}-V_{L}W_{R}^{3}+V_{L}^{4}+W_{R}^{4}\right).$
-The factor $\left(V_{L}^{5}W_{L}^{5}-V_{R}^{5}W_{R}^{5}\right)$ is also present for the non-zero field case, and is solved by (B.3), however it does not lead to exceptional points.
+$\times\left(-V_{L}^{3}W_{R}+V_{L}^{2}W_{R}^{2}-V_{L}W_{R}^{3}+V_{L}^{4}+W_{R}^{4}\right).$
+The factor $\left(V_{L}^{5}W_{L}^{5}-V_{R}^{5}W_{R}^{5}\right)$ is also present for the non-zero field case, and is solved by (A.4), however it does not lead to exceptional points.
 To analyze the factor $\left(\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}\right)U(u)$ we recall (III.25) for every possible quantized $k$ (five for $N=10$).
 For example, for $k=2\pi$,
-$H_{1}(2\pi)H_{2}(2\pi)-u^{2}=\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}.$ (B.6)
+$H_{1}(2\pi)H_{2}(2\pi)-u^{2}=\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}.$ (A.7)
 Actually, we can check by brute force the following factorization,
 $\left(\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}\right)U(u)=$
-$\prod_{m=1}^{5}H_{1}(k_{m})H_{2}(k_{m})-u^{2}$ (B.7)
+$\prod_{m=1}^{5}H_{1}(k_{m})H_{2}(k_{m})-u^{2}$ (A.8)
 and confirm that the exceptional points give the roots of the discriminant.
-B.2 Open boundary conditions
-Now let us consider $N=11$, $\gamma=0$ (OBC) and $u=0$. In this case the discriminant (B.1) is given by,
-$\text{disc}(P_{11}^{\text{(OBC)}}(z))=48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}f_{1}^{3}f_{2}^{3}f_{3}^{3}$ (B.8)
+A.2 Open boundary conditions
+Now let us consider $N=11$, $\gamma=0$ (OBC) and $u=0$. In this case the discriminant (A.1) is given by,
+$\text{disc}(P_{11}^{\text{(OBC)}}(z))=48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}f_{1}^{3}f_{2}^{3}f_{3}^{3}$ (A.9)
 where
@@
 $f_{2}$ $=$ $V_{L}^{2}V_{R}^{2}-V_{L}V_{R}W_{L}W_{R}+W_{L}^{2}W_{R}^{2}$
-$f_{3}$ $=$ $V_{L}^{2}V_{R}^{2}+V_{L}V_{R}W_{L}W_{R}+W_{L}^{2}W_{R}^{2}.$ (B.9)
-Of course, the discriminant (B.8) vanishes when any of the couplings vanishes, consistent with our findings in the main text. The remaining factors can be verified by rewriting
-$f_{1}f_{2}f_{3}=V_{L}^{5}V_{R}^{5}\sum_{j=0}^{5}\left(\frac{1}{\sqrt{\delta}}\right)^{2j}$ (B.10)
+$f_{3}$ $=$ $V_{L}^{2}V_{R}^{2}+V_{L}V_{R}W_{L}W_{R}+W_{L}^{2}W_{R}^{2}.$ (A.10)
+Of course, the discriminant (A.9) vanishes when any of the couplings vanishes, consistent with our findings in the main text. The remaining factors can be verified by rewriting
+$f_{1}f_{2}f_{3}=V_{L}^{5}V_{R}^{5}\sum_{j=0}^{5}\left(\frac{1}{\sqrt{\delta}}\right)^{2j}$ (A.11)
 whose ten roots are given by (IV.16).
 Next, we turn on the field $u$. In this case, we have,
-$\text{disc}(P_{11}^{\text{(OBC)}}(z))$ $=$ $-48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}$ (B.11)
+$\text{disc}(P_{11}^{\text{(OBC)}}(z))$ $=$ $-48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}$ (A.12)
 $\times$ $f_{1}^{2}f_{2}^{2}f_{3}^{2}g_{1}g_{2}g_{3}$
-where $f_{j}$ is given by (B.2) and $g_{j}$ are given by
+where $f_{j}$ is given by (A.2) and $g_{j}$ are given by
 $g_{1}$ $=$ $u^{2}-(V_{L}V_{R}+W_{L}W_{R}),$
 $g_{2}$ $=$ $u^{4}-2u^{2}f_{1}+f_{2},$
-$g_{3}$ $=$ $u^{4}-2u^{2}f_{1}+f_{3}.$ (B.12)
+$g_{3}$ $=$ $u^{4}-2u^{2}f_{1}+f_{3}.$ (A.13)
 Part of the roots remain the same as in the zero field case, but field dependent factors $g_{j}$ appear. These are associated with (IV.20). Indeed, by brute force computation we can verify
-$\prod_{m=1}^{5}u^{2}-\tilde{H}_{1}(\theta_{m})\tilde{H}_{2}(\theta_{m})=g_{1}g_{2}g_{3}$ (B.13)
+$\prod_{m=1}^{5}u^{2}-\tilde{H}_{1}(\theta_{m})\tilde{H}_{2}(\theta_{m})=g_{1}g_{2}g_{3}$ (A.14)
 where $\theta_{m}$ is given by (IV.8). Let us remark that the product $g_{1}g_{2}g_{3}$ only depends on $\bar{u}$ and $\sqrt{\delta}$, up to an overall factor.
 Finally, let us consider $N=12$ with non-zero field $u$. The discriminant is given by,
-$\text{disc}(P_{12}^{\text{(OBC)}}(z))=4096V_{L}^{20}V_{R}^{20}W_{L}^{30}W_{R}^{30}U_{0}U_{1}(u)$ (B.14)
+$\text{disc}(P_{12}^{\text{(OBC)}}(z))=4096V_{L}^{20}V_{R}^{20}W_{L}^{30}W_{R}^{30}U_{0}U_{1}(u)$ (A.15)
 where
-$U_{0}$ $=$ $153664V_{L}^{5}V_{R}^{5}+98000V_{L}^{4}V_{R}^{4}W_{L}W_{R}$ (B.15)
+$U_{0}$ $=$ $153664V_{L}^{5}V_{R}^{5}+98000V_{L}^{4}V_{R}^{4}W_{L}W_{R}$ (A.16)
 $+$ $60152V_{L}^{3}V_{R}^{3}W_{L}^{2}W_{R}^{2}+34709V_{L}^{2}V_{R}^{2}W_{L}^{3}W_{R}^{3}$
@@
 and
-$U_{1}(u)$ $=$ $u^{12}-u^{10}(6V_{L}V_{R}+5W_{L}W_{R})$ (B.16)
+$U_{1}(u)$ $=$ $u^{12}-u^{10}(6V_{L}V_{R}+5W_{L}W_{R})$ (A.17)
 $+$ $5u^{8}(3V_{L}^{2}V_{R}^{2}+4V_{L}V_{R}W_{L}W_{R}+2W_{L}^{2}W_{R}^{2})$
@@
 $+$ $V_{L}^{6}V_{R}^{6}.$
-Again, it is clear that the vanishing of any of the coupling parameters leads to a null discriminant (B.14). The factor $U_{0}$ can be written as
+Again, it is clear that the vanishing of any of the coupling parameters leads to a null discriminant (A.15). The factor $U_{0}$ can be written as
 $U_{0}=153664V_{L}^{5}V_{R}^{5}\Big(1+\frac{125}{196}\frac{1}{\sqrt{\delta}^{2}}+\frac{7519}{19208}\frac{1}{\sqrt{\delta}^{4}}$
-$+\frac{34709}{153664}\frac{1}{\sqrt{\delta}^{6}}+\frac{279}{2401}\frac{1}{\sqrt{\delta}^{8}}+\frac{108}{2041}\frac{1}{\sqrt{\delta}^{10}}\Big{)}.$ (B.17)
+$+\frac{34709}{153664}\frac{1}{\sqrt{\delta}^{6}}+\frac{279}{2401}\frac{1}{\sqrt{\delta}^{8}}+\frac{108}{2041}\frac{1}{\sqrt{\delta}^{10}}\Big).$ (A.18)
 This factor is independent of the field $u$, and we can check numerically that the ten roots are given by (IV.25,IV.26), see summary in Table 1.
@@
 $-\frac{\bar{u}(\bar{u}-1)^{2}(\bar{u}+1)^{3}}{\sqrt{\delta}^{6}}\Bigg)\Bigg(1-\frac{\bar{u}(3\bar{u}-1)}{\sqrt{\delta}^{2}}$
-$+\frac{\bar{u}(\bar{u}-1)(3\bar{u}^{2}+\bar{u}-1)}{\sqrt{\delta}^{4}}-\frac{\bar{u}(\bar{u}+1)^{2}(\bar{u}-1)^{3}}{\sqrt{\delta}^{6}}\Bigg{)}.$ (B.18)
+$+\frac{\bar{u}(\bar{u}-1)(3\bar{u}^{2}+\bar{u}-1)}{\sqrt{\delta}^{4}}-\frac{\bar{u}(\bar{u}+1)^{2}(\bar{u}-1)^{3}}{\sqrt{\delta}^{6}}\Bigg).$ (A.19)
 We verify numerically that (IV.22,IV.23) do provide the roots of $U_{1}(u)=0$.
 In Table 2 we consider $\bar{u}=1/2$ and the associated solutions of (IV.22), as well as the associated exceptional $\sqrt{\delta}$’s.
-Table 1: The ten roots of Eq. (B.2).
+Table 1: The ten roots of Eq. (A.2).
 $\theta_{\text{EP}}^{\prime}$ $\sqrt{\delta}$
-$05$ 0.209455 i $-0.628308+0.408561\text{i}$
-$0.580505+0.209455\text{i}$ $-0.628308-0.408561\text{i}$
-$1.07762-0.24362\text{i}$ $-0.344178+0.638086\text{i}$
-$1.07762+0.24362\text{i}$ $-0.344178-0.638086\text{i}$
+$0.580505-0.209455\text{i}$ $0.628308+0.408561\text{i}$
+$0.580505+0.209455\text{i}$ $0.628308-0.408561\text{i}$
+$1.07762-0.24362\text{i}$ $0.344178+0.638086\text{i}$
+$1.07762+0.24362\text{i}$ $0.344178-0.638086\text{i}$
 $1.5708-0.252951\text{i}$ $0.718376$
-$1.5708+0.252951\text{i}$ $-0.718376$
+$1.5708+0.252951\text{i}$ 0.718376-0.718376
 $2.06398-0.24362\text{i}$ $0.344178+0.638086\text{i}$
@@
 $2.56109+0.209455\text{i}$ $0.628308-0.408561\text{i}$
-Table 2: We consider $\bar{u}=0.5$ First six lines, $+$ solutions, last six lines -- solutions.
+Table 2: We consider $\bar{u}=0.5$ First six lines, ++ solutions, last six lines - solutions.
 $\theta_{\text{EP}}^{\prime}$ $\sqrt{\delta}$
-$0.642557-0.0955613\text{i}$ $-0.673088+0.294372\text{i}$
-$0.642557+0.0955613\text{i}$ $-0.673088-0.294372\text{i}$
-[…]
-$1.5708+0.191123\text{i}$ $-0.694828\text{i}$
+$0.642557-0.0955613\text{i}$ $0.673088+0.294372\text{i}$
+$0.642557+0.0955613\text{i}$ $0.673088-0.294372\text{i}$
+$1.5708-0.191123\text{i}$ $0.694828\text{i}$
+$1.5708+0.191123\text{i}$ $0.694828\text{i}$
 $2.49904-0.0955613\text{i}$ $0.673088+0.294372\text{i}$
 $2.49904+0.0955613\text{i}$ $0.673088-0.294372\text{i}$
-$0.343297$ $-1.31139$
-$1.10745-0.172812\text{i}$ $-0.36085+0.604218\text{i}$
-$1.10745+0.172812\text{i}$ $-0.36085-0.604218\text{i}$
+$0.343297$ 1.31139-1.31139
+$1.10745-0.172812\text{i}$ $0.36085+0.604218\text{i}$
+$1.10745+0.172812\text{i}$ $0.36085-0.604218\text{i}$
 $2.03414-0.172812\text{i}$ $0.36085+0.604218\text{i}$
@@
 $2.7983$ $1.31139$
+Appendix B Some details for the open boundary case
+To make this paper self-contained, we review some details of the diagonalization of $\mathcal{T}_{N}$ (II) for open boundary conditions, that is, $\gamma=0$, following [49, 50].
+The approach in [49] is based on recurrence relations satisfied by the characteristic polynomial,
+$P_{N}(z)=\det_{N}\left(\mathcal{T}_{N}-z\right).$ (B.1)
+The roots $P_{N}(z)=0$ are the quasienergies of the Hamiltonian (II.1). The characteristic polynomial (B.1) for $\gamma=0$ is special as it satisfies the recurrence relations
+$P_{2j}(z)$ $=$ $(-\text{i}u-z)P_{2j-1}(z)-V_{L}V_{R}P_{2j-2}(z),$
+$P_{2j-1}(z)$ $=$ $(\text{i}u-z)P_{2j-2}(z)-W_{L}W_{R}P_{2j-3}(z)$ (B.2)
+depending on the parity of $N$. Note that one has an additional recurrence relation for the characteristic polynomial, namely
+$P_{j}(z)$ $=$ $\left((z+\text{i}u)(z-\text{i}u)-(V_{L}V_{R}+W_{L}W_{R})\right)P_{j-2}(z)$ (B.3)
+- $V_{L}V_{R}W_{L}W_{R}P_{j-4}(z),$
+where appropriate initial conditions must be chosen.
+On the other hand, following [50], one notes that equations (II.7,II.8) can be iterated to the same fourth order difference equation,
+$r_{j+2}+\frac{V_{R}W_{R}}{V_{L}W_{L}}r_{j-2}$
+$=\left(-\frac{(\text{i}u+\epsilon)(\text{i}u-\epsilon)+(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}}\right)r_{j}.$ (B.4)
+The difference equation (B) can be compared to (B.3) leading to the solution (IV.10).
+In addition, following [50], introduce
+$T_{n}(\alpha,\beta)=\frac{\alpha^{n+1}-\beta^{n+1}}{\alpha-\beta}$ (B.5)
+and the sequence
+$a_{n+2}-(\alpha+\beta)a_{n+1}+\alpha\beta a_{n}=0.$ (B.6)
+Then, it follows that
+$a_{n}=T_{n-2}(\alpha,\beta)a_{2}-\alpha\beta T_{n-3}(\alpha,\beta)a_{1}$ (B.7)
+is the solution of the difference equation (B.6). Now choosing
+$\alpha+\beta=\frac{(\text{i}u-\epsilon)(-\text{i}u-\epsilon)-(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}},$ (B.8)
+$\alpha\beta=\frac{V_{R}W_{R}}{V_{L}W_{L}},$ (B.9)
+it follows that
+$r_{2j}=T_{j-2}(\alpha,\beta)r_{4}-\alpha\beta T_{j-3}(\alpha,\beta)r_{2},$ (B.10)
+$r_{2j-1}=T_{j-2}(\alpha,\beta)r_{3}-\alpha\beta T_{j-3}(\alpha,\beta)r_{1},$ (B.11)
+is a solution of (II.7,II.8,B). The boundary conditions determine,
+$r_{1}$ $=$ $V_{L},\quad r_{2}=\epsilon-\text{i}u,\quad r_{3}=W_{R}+V_{L}(\alpha+\beta),$
+$r_{4}$ $=$ $(\epsilon-\text{i}u)\left(2\frac{W_{R}}{V_{L}}+\alpha+\beta\right).$ (B.12)
+where $r_{1}=V_{L}$ is an arbitrary choice. The conditions (B.8) give $\{\epsilon,\beta\}$ in terms of $\alpha$,
+$\epsilon$ $=$ $\pm\sqrt{V_{L}V_{R}+W_{L}W_{R}+V_{L}W_{L}(\alpha+\beta)-u^{2}},$
+$\beta$ $=$ $\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}.$ (B.13)
+Using the parametrization
+$\alpha=\frac{\sqrt{V_{R}}\sqrt{W_{R}}}{\sqrt{V_{L}}\sqrt{W_{L}}}\exp(\text{i}\theta),$ (B.14)
+and then
+$T_{n}\left(\alpha,\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}\right)=\frac{V_{R}^{\frac{n}{2}}W_{R}^{\frac{n}{2}}}{V_{L}^{\frac{n}{2}}W_{L}^{\frac{n}{2}}}\frac{\sin((n+1)\theta)}{\sin(\theta)}$ (B.15)
+in (B.10,B.11) we find (IV.10,LABEL:vLodd) with the initial conditions (B).
+The boundary condition at the end of the chain gives a condition on the parameter $\theta$, which depends on whether $N$ is odd or even.
+For odd $N$, the condition $x_{L+1}=0$ comes from (IV.10) with $2n=L+1$ and gives the quantized solution (IV.8).
+For even $N$, the condition $x_{L+1}=0$ comes from (LABEL:vLodd) with 2n1=L+12n-1=L+1 and leads to the transcendental equation (IV.9).
+The solution (IV.12,LABEL:vLoddL) is obtained analogously.